### v2

# Global well-posedness for intermediate NLS with nonvanishing conditions at infinity

Abstract.
The intermediate nonlinear Schrödinger equation models quasi-harmonic internal waves in two-fluid layer system, and admits dark solitons, that is, solutions with nonvanishing boundary conditions at spatial infinity. These solutions fall outside existing well-posedness theories. We establish local and global well-posedness in a Zhidkov-type space naturally suited to such non-trivial boundary conditions, and extend these results to a generalized defocusing equation. This appears to be the first well-posedness result for the equation in a functional setting adapted to its dark soliton structure.
Key words and phrases:
Derivative nonlinear Schrödinger equation, integrable systems, intermediate NLS, Calogero-Moser, global well-posedness, modified energy method, Zhidkov space.
1. Introduction
We study the intermediate nonlinear Schrödinger equation in one space dimension,
(INLS) $i\partial_{t}u-\partial_{x}^{2}u=u(i-T_{\delta})\partial_{x}|u|^{2}\qquad\mbox{in}~\mathbb{R}\times\mathbb{R},$
where $u$ is a complex-valued function, $\delta>0$ is a parameter called the total fluid depth, and $T_{\delta}$ is a singular integral operator defined by
$T_{\delta}(f)(x):=\frac{1}{2\delta}\lim_{\varepsilon\downarrow 0}\int_{|x-y|>\varepsilon}\coth{\Big(\frac{\pi(x-y)}{2\delta}\Big)}f(y)\,dy.$
The equation (INLS) was introduced by Pelinovsky in [39] to describe the long-term dynamics of the envelope of weakly nonlinear internal waves in a stratified fluid of finite depth.
The operator $T_{\delta}$ can be written as a Fourier multiplier form111 We will give a proof of the formula (1.1) in Appendix A.
(1.1) $\widehat{T_{\delta}(f)}(\xi)=-i\coth(\delta\xi)\hat{f}(\xi),$
where $\widehat{\cdot}$ denotes the Fourier transformation (see Section 2.1). The asymptotics of this symbol reveal that (INLS) interpolates between two fundamental models. In the deep-water limit $\delta\to\infty$, the symbol $i\coth(\delta\xi)$ converges to $i\,\mathrm{sgn}(\xi)$, so that $T_{\delta}$ converges (at least formally) to the Hilbert transform $H$, where
$H(f)(x):=\lim_{\varepsilon\downarrow 0}\int_{|x-y|>\varepsilon}\frac{f(y)}{\pi(x-y)}\,dy,\qquad\widehat{H(f)}(\xi)=-i\,\mathrm{sgn}(\xi)\hat{f}(\xi).$
Thus, as shown in [9, Theorem 1.3], when $\delta\to\infty$, (INLS) reduces to the defocusing Calogero-Moser derivative nonlinear Schrödinger equation
(CM-defoc) $i\partial_{t}u-\partial_{x}^{2}u=u\,(i-H)\partial_{x}|u|^{2}$
originally derived on the real line in [39, 30, 19], and extended to the periodic setting in [34, 4, 6]. The equation (CM-defoc) has been deeply studied in recent years, where global well-posedness has been derived in the Hardy space $L^{2}_{+}(\mathbb{R})$ in [25] and in $L^{2}_{+}(\mathbb{T})$ in [4]. The focusing variant
(CM-foc) $i\partial_{t}u-\partial_{x}^{2}u=u\,(i+H)\partial_{x}|u|^{2}$
has also been shown to be globally well-posed in $L^{2}_{+}(\mathbb{R})$ [19, 25] for initial data $\|u(0)\|_{L^{2}}^{2}<2\pi$ and in $L^{2}_{+}(\mathbb{T})$ for small initial data [4], with the bound $2\pi$ known to be optimal thanks to the work of [22, 26, 27, 24]. Recently, Chapouto–Forlano–Laurens [9] proved that the well-posedness theory for (CM-foc) extends beyond the Hardy-space setting. They established local well-posedness in $H^{s}(\mathbb{R})$ for all $s>1/4$, and global well-posedness for initial data with small $L^{2}$-norm. Solitons for (CM-foc) were derived in [19, 30, 31, 36, 33] on the real line, and they display turbulent behavior [19]. On the torus, the set of special solutions, including traveling waves, is much wider than on $\mathbb{R}$ [6, 30, 36, 33]. We also mention that the zero dispersion limit on $\mathbb{R}$ was studied in [34, 5], and the inverse scattering theory for the Calogero-Moser model in the defocusing [35] and focusing [17, 10] cases has also been investigated.
On the other hand, in the shallow-water limit $\delta\to 0$, after the rescaling $v(x,t)=\sqrt{\delta}u(x,t)$, (INLS) formally converges to the defocusing nonlinear Schrödinger equation
(NLS) $i\partial_{t}v-\partial_{x}^{2}v=-v|v|^{2}.$
This equation and the focusing variant are famous integrable equations solved by the method of inverse scattering transformation [1, 43, 16, 29, 20]. Recently, Harrop-Griffiths–Killip–Vişan [21] proved that (NLS) and the focusing variant are globally well-posed in $H^{s}(\mathbb{R})$ for any s>1/2s>-1/2. This result is sharp in the sense that the norm inflation occurs in $H^{s}(\mathbb{R})$ for $s\leq-1/2$ (see, e.g., [12, 28]). See, e.g., [41, 21] for the development of well-posedness of nonlinear Schrödinger equations.
A notable property of (INLS) is that it admits dark solitons [39, 37, 32]: for any $\rho>0$ and integer $N\geq 1$, there exists an $N$-soliton solution $Q$ which satisfies (INLS) on the whole time interval $\mathbb{R}$ and a nonvanishing boundary condition $|Q(x,t)|\to\rho$ as $|x|\to\infty$ for all $t\in\mathbb{R}$, while $\partial_{x}Q$ remains in $L^{2}(\mathbb{R})$. Thus, the existence of dark solitons naturally suggests that the appropriate functional framework for analyzing such solutions to (INLS) should involve nonvanishing boundary conditions. Accordingly, we introduce the following subsets of Zhidkov space:
(1.2) $\begin{gathered}\mathcal{Z}^{k}_{\rho}:=\{u\in L^{\infty}(\mathbb{R}):\partial_{x}u\in H^{k-1}(\mathbb{R}),\,|u|^{2}-\rho^{2}\in L^{2}(\mathbb{R})\}\qquad(\rho\geq 0),\\ \mathcal{Z}^{k}:=\bigcup_{\rho\geq 0}\mathcal{Z}^{k}_{\rho}=\{u\in L^{\infty}(\mathbb{R}):\partial_{x}u\in H^{k-1}(\mathbb{R}),\,|u|^{2}-\rho^{2}\in L^{2}(\mathbb{R})~\text{for some $\rho\geq 0$}\}.\end{gathered}$
This framework naturally accommodates dark soliton solutions, which are localized deviation from a constant background. We will use the spaces $\mathcal{Z}^{k}$ only for positive integers $k$. Note that this is not a vector space but a complete metric space with the following distance function:
(1.3) $d^{k}(f,g):=\|\partial_{x}(f-g)\|_{H^{k-1}}+\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}},$
where for $f\in\mathcal{Z}^{k}$ we denote by $\rho_{f}$ the unique non-negative number such that $|f|^{2}-\rho_{f}^{2}\in L^{2}(\mathbb{R})$. We also note that for each $\rho\geq 0$, $\mathcal{Z}^{k}_{\rho}$ is a closed subset of $\mathcal{Z}^{k}$, and that $H^{k}\hookrightarrow\mathcal{Z}^{k}_{0}$. To measure the size of functions in $\mathcal{Z}^{k}$, we introduce the functional
(1.4) $\mathcal{E}^{k}(u):=\|\partial_{x}u\|_{H^{k-1}}^{2}+\|u\|_{L^{\infty}}^{2}+\big\||u|^{2}-\rho_{u}^{2}\big\|_{L^{2}}.$
Although the existence of dark solitons for (INLS) has been extensively studied in the literature [39, 37, 32, 30], and their asymptotic behavior has been investigated in [2], the problem of local well-posedness in a functional setting appropriate for these solitons remained open. In [15], local well-posedness222 Actually, he proved the local well-posedness of a family of nonlocal nonlinear Schrödinger equation [38] that includes also (INLS). was established in the Sobolev space $H^{s}(\mathbb{R})$ for $s\geq 1$. Later, Barros–de Moura–Santos have shown in [7] the local well-posedness of (INLS) for small initial data in the Besov space $B_{2}^{1/2,1}(\mathbb{R})$. More recently, Chapouto–Forlano–Laurens have proven in [9] that (INLS) is globally well-posed in $H^{s}(\mathbb{R})$ with $s>1/4$ for small $L^{2}$ initial data. However, this setting does not accommodate the nonvanishing boundary conditions intrinsic to dark solitons. The aim of the present work is therefore to address this gap by proving that (INLS) is globally well-posed in the space $\mathcal{Z}^{2}$. One could likely employ weighted Sobolev spaces to accommodate dark solitons; however, this approach lies beyond the scope of the present paper.
Our argument does not rely on the complete integrability of (INLS). Thus, we can deal with the following generalized version of (INLS):
(gINLS) $i\partial_{t}u-\partial_{x}^{2}u=u(i\alpha-\beta T_{\delta})\partial_{x}|u|^{2}\qquad\text{in $\mathbb{R}\times\mathbb{R}$,}$
where $\alpha,\beta\in\mathbb{R}$, and $u$ is a complex-valued function. Note that (gINLS) becomes completely integrable when $\beta=\pm|\alpha|\neq 0$, with the plus (resp. minus) sign corresponding to the defocusing (resp. focusing) nonlinearity.333 The sign of $\alpha$ is not relevant. Indeed, if $\alpha\neq 0$ one can set $\alpha=1$ by the change of variables $u(x,t)\mapsto|\alpha|^{1/2}u((\operatorname{sgn}{\alpha})x,t)$, by which the nonlinearity is transformed into $u(i-\beta|\alpha|^{-1}T_{\delta})\partial_{x}|u|^{2}$. The original (INLS) can be recovered by setting $\alpha=\beta=1$.
Our first main result establishes local well-posedness in $\mathcal{Z}^{2}$ for (gINLS).
Theorem 1.1 (Local well-posedness in $\mathcal{Z}^{2}$).
Let $\delta>0$ and $\alpha,\beta\in\mathbb{R}$. For any $M_{0}>0$, there exists $T_{*}\in(0,1]$ satisfying
$T_{*}\geq c_{\alpha,\beta}\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$
such that for any initial datum $\phi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$, the following hold:
  • There exists a unique solution $u\in C([0,T_{*}];\mathcal{Z}^{2})$ to (gINLS) with $u(0)=\phi$. Moreover, the solution map $\phi\mapsto u$ is continuous from $\{\phi\in\mathcal{Z}^{2}\colon\mathcal{E}^{2}(\phi)\leq M_{0}^{2}\}$ into $C([0,T_{*}];\mathcal{Z}^{2})$.
  • The solution preserves the background density: $u\in C([0,T_{*}];\mathcal{Z}^{2}_{\rho_{\phi}})$.
  • (Persistence of regularity) If $\phi\in\mathcal{Z}^{k}$ for some integer $k\geq 3$, then $u\in C([0,T_{*}];\mathcal{Z}^{k})$.
Remark 1.1.
Notice that the $L^{2}$-rescaling
$u_{\lambda}(x,t)=\lambda^{-\frac{1}{2}}u(\lambda^{-1}x,\lambda^{-2}t)$
maps a solution $u$ of (INLS) to another solution of (INLS), but with rescaled depth $\lambda\delta$. This formally indicates that $L^{2}(\mathbb{R})$ is the scaling-critical space. Consequently, one may expect well-posedness for initial data in $\mathcal{Z}^{0}$, a problem that remains widely open, particularly in the non-decaying data setting.
The proof relies on a modified energy method [40, 13] enabling us to control the derivative present in the nonlinearity of (gINLS), and to avoid the direct use of a standard fixed-point argument in the space $\mathcal{Z}^{2}$. We remark that in [11], using Kato’s iterative scheme similar to [14], Chen claimed a well-posedness result in the space $\mathcal{Z}^{2}_{\rho}$ with $\rho=1$ for the defocusing Calogero-Moser DNLS equation (CM-defoc). Observe that the existence time $T_{*}$ in Theorem 1.1 is uniformly bounded from below when $\delta\to\infty$. In fact, our proof is easily adapted to the limiting case $\delta=\infty$, yielding well-posedness in $\mathcal{Z}^{2}$ for the following generalized version of the Calogero-Moser DNLS equation:
(gCM) $i\partial_{t}u-\partial_{x}^{2}u=u(i\alpha-\beta H)\partial_{x}|u|^{2}\qquad\text{in $\mathbb{R}\times\mathbb{R}$}.$
Corollary 1.2 (Local well-posedness for (gCM)).
Let $\alpha,\beta\in\mathbb{R}$ and $M_{0}>0$. The same local well-posedness result as in Theorem 1.1 holds for (gCM), with $T_{*}\in(0,1]$ satisfying
$T_{*}\geq c_{\alpha,\beta}(1+M_{0}^{4})^{-1}.$
Beyond the independent well-posedness of both systems, the following result establishes the rigorous convergence of the (gINLS) flow to the (gCM) flow in the deep-water limit $\delta\to\infty$ for solutions with nonvanishing boundary conditions.
Corollary 1.3 (Convergence in the deep-water limit).
Let $\alpha,\beta\in\mathbb{R}$ be arbitrary parameters. Suppose that $\{\phi_{\delta}\}_{\delta\in[1,\infty]}\subset\mathcal{Z}^{2}$ is a family of initial data satisfying
$\lim_{\delta\to\infty}d^{2}(\phi_{\delta},\phi_{\infty})=0.$
Let $u_{\delta}$ for $\delta\in[1,\infty)$ and $u_{\infty}$ denote the unique maximal solutions in $C([0,T_{\delta,\max});\mathcal{Z}^{2})$ to the Cauchy problems for (gINLS) with $u_{\delta}(0)=\phi_{\delta}$ and (gCM) with $u_{\infty}(0)=\phi_{\infty}$, respectively, as constructed in Theorem 1.1 and Corollary 1.2. Then, the maximal existence time $T_{\delta,\max}$ satisfies the lower semicontinuity property:
$\liminf_{\delta\to\infty}T_{\delta,\max}\geq T_{\infty,\max}.$
Furthermore, for any $0<T<T_{\infty,\max}$, the solutions converge in the sense that
$\lim_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T}}=0.$
For the intermediate model (INLS), the analysis is more subtle than (NLS) and (CM-defoc) because of the singular nature of the operator $T_{\delta}$. Although (INLS) admits several Lax representations and infinite number of conservation laws [37, 9], these integrable structures do not provide well-defined conserved quantities in a setting with nonvanishing boundary conditions such as $\mathcal{Z}^{k}$. Indeed, if we consider the conservation laws at the $H^{n/2}$ level, derived via the Lax operator given in [9]:
(1.5) $\int_{\mathbb{R}}\overline{u}\mathcal{L}_{u}^{n}u\,dx\quad(n=0,1,2,\dots),\qquad\mathcal{L}_{u}:\ f\in H^{1}(\mathbb{R})\mapsto-i\partial_{x}f+u\frac{1+iT_{\delta}}{2}(\overline{u}f)\in L^{2}(\mathbb{R}),$
then, after expansion and reduction of the number of $T_{\delta}$ operators using Cotlar’s identity (see [9, Eq. (2.12)]), the first five conservation laws read
$E_{0}(u)$ $=\int_{\mathbb{R}}|u|^{2}\,dx,$
$E_{1/2}(u)$ $=\int_{\mathbb{R}}\Big\{\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{1}{2}|u|^{4}\Big\}\,dx,$
$E_{1}(u)$ $=\int_{\mathbb{R}}\Big\{|\partial_{x}u|^{2}+|u|^{2}\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{1}{2}|u|^{2}T_{\delta}\partial_{x}(|u|^{2})+\frac{1}{3}|u|^{6}\Big\}\,dx,$
$E_{3/2}(u)$ $=\int_{\mathbb{R}}\Big\{\begin{aligned} &\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]+\frac{1}{2}[\partial_{x}(|u|^{2})]^{2}+\frac{3}{2}|u|^{2}|\partial_{x}u|^{2}+\frac{3}{2}\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}(|u|^{2})\\ &+|u|^{4}\Big(\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{3}{4}T_{\delta}\partial_{x}(|u|^{2})\Big)+\frac{1}{4}|u|^{8}\Big\}\,dx,\end{aligned}$
$E_{2}(u)$ $=\int_{\mathbb{R}}\Big\{\begin{aligned} &|\partial_{x}^{2}u|^{2}+\Big(2|\partial_{x}u|^{2}-3\partial_{x}^{2}(|u|^{2})\Big)\operatorname{Im}[\overline{u}\partial_{x}u]\\ &+\Big(2|\partial_{x}u|^{2}-\frac{3}{4}\partial_{x}^{2}(|u|^{2})\Big)T_{\delta}\partial_{x}(|u|^{2})+\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\\ &+2|u|^{4}|\partial_{x}u|^{2}+|u|^{2}\Big(\frac{3}{2}[\partial_{x}(|u|^{2})]^{2}+\frac{1}{2}[T_{\delta}\partial_{x}(|u|^{2})]^{2}\Big)\\ &+\operatorname{Im}[\overline{u}\partial_{x}u]\Big(2|u|^{2}T_{\delta}\partial_{x}(|u|^{2})+T_{\delta}\partial_{x}(|u|^{4})\Big)\\ &+|u|^{6}\Big(\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{2}{3}T_{\delta}\partial_{x}(|u|^{2})\Big)+\frac{1}{4}|u|^{4}T_{\delta}\partial_{x}(|u|^{4})+\frac{1}{5}|u|^{10}\Big\}\,dx.\end{aligned}$
However, these integrals fail to be well-defined for $u\in\mathcal{Z}^{2}$ with $\rho_{u}>0$, even after substituting $|u|^{2(n+1)}-\rho_{u}^{2(n+1)}$ for the last term $|u|^{2(n+1)}$ in $E_{n/2}(u)$. To overcome this, we construct conservation laws via the modified energy method, which allows us to establish global well-posedness for (gINLS). Our motivation for considering the more general model (gINLS), rather than (INLS), is to emphasize that the argument does not rely on the complete integrability of (INLS).
Theorem 1.4 (Global well-posedness and a priori bounds).
Let $\delta>0$. For general $\alpha\neq 0$ and $\beta\geq 0$, the Cauchy problem is globally well-posed in
$\mathcal{Z}^{2}_{*}:=\bigcup_{\rho>0}\mathcal{Z}^{2}_{\rho}$
for both (gINLS) and (gCM). Furthermore, the energy functional satisfies a uniform bound in the integrable–defocusing cases:
$\sup_{t\in\mathbb{R}}\mathcal{E}^{2}\big(u(t)\big)<\infty,$
which holds for (INLS) as well as for (CM-defoc).
Remark 1.2.
In the general defocusing case $\alpha\neq 0$, $\beta\geq 0$, we still have a uniform bound for $\mathcal{E}^{1}(u(t))$ but only an exponential bound for $\|\partial_{x}^{2}u(t)\|_{L^{2}}$; see Theorem 4.1 in Section 4. The global theory is built on two new conservation laws compatible with nonvanishing boundary conditions, which allow us to control $\|u\|_{L^{\infty}}^{2},\|\partial_{x}u\|_{H^{1}}^{2},$ and $\||u|^{2}-\rho_{u}^{2}\|_{L^{2}}$ in time.
The paper is organized as follows. Section 2 collects preliminary estimates and properties of $T_{\delta}$ and $\mathcal{Z}^{k}$. Section 3 proves local well-posedness in $\mathcal{Z}^{2}$. Section 4 establishes the modified conservation laws and global well-posedness in $\mathcal{Z}^{2}_{*}$.
Acknowledgments. T.A. was supported by JSPS KAKENHI Grants Number 20K03697 and 25K07082. R.B. was supported by an AMS-Simons Travel Grant. R.B. gratefully acknowledges the University of Victoria and the Pacific Institute for the Mathematical Sciences (PIMS) for their support of the visit to UVic. S.I. acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (NSERC) under Grant No. 371637-2025. Part of this work was carried out during an extended visit by S.I. to the Research Institute for Mathematical Sciences (RIMS). S.I. is grateful to colleagues and the staff at RIMS for their warm hospitality and stimulating research environment. Part of this work was carried out during visits by R.B. and S.I. to Shizuoka University. They would like to thank their colleagues and the staff at Shizuoka University for their warm hospitality. N.K. was supported by JSPS KAKENHI Grant Number 25K07066.
2. Preliminaries
We begin by recording the notation used throughout the paper and present several useful inequalities that will be invoked repeatedly in the analysis.
2.1. Notation
We use $\mathcal{F}$ (or $\widehat{\,\cdot\,}$) and $\mathcal{F}^{-1}$ to denote the Fourier and its inverse transformations:
$\mathcal{F}[f](\xi)=\widehat{f}(\xi):=\int_{\mathbb{R}}e^{-ix\xi}f(x)\,dx,\qquad\mathcal{F}^{-1}[g](x):=\frac{1}{2\pi}\int_{\mathbb{R}}e^{ix\xi}g(\xi)\,d\xi.$
We use $U(t)$ to denote the free Schrödinger operator given by
$U(t):=e^{-it\partial_{x}^{2}}=\mathcal{F}^{-1}e^{it|\xi|^{2}}\mathcal{F}.$
Our convention for the $L^{2}$-inner product $\left(f,g\right)_{L^{2}}$ is
$\left(f,g\right)_{L^{2}}:=\int_{\mathbb{R}}f(x)\overline{g}(x)\,dx.$
For the frequency localization, let $\varphi$ be a smooth, nonnegative radial function supported on $|\xi|\leq 2$ with $\varphi(\xi)=1$ for $|\xi|\leq 1$. For each nonnegative integer $j\in\mathbb{Z}_{\geq 0}$, define the (inhomogeneous) Littlewood-Paley projection $P_{\leq j}$ as the Fourier multiplier with symbol $\varphi(\xi/2^{j})$, and set $P_{\geq j+1}:=I-P_{\leq j}$, $P_{j}:=P_{\leq j}-P_{\leq j-1}$ ($j\geq 1$) and $P_{0}:=P_{\leq 0}$.
For $T>0$, we use the convention $L_{T}^{\infty}$ to denote the space $L^{\infty}(0,T)$. Furthermore, for a Banach space $B$, we use $L^{\infty}_{T}B$ to denote $L^{\infty}(0,T;B)$.
We use $\mathcal{N}(u)$ to denote the nonlinear term of (gINLS), namely
$\mathcal{N}(u):=iu(\alpha+i\beta T_{\delta})\partial_{x}|u|^{2},$
and also define the operator $\mathcal{L}_{\delta}$ as
$\mathcal{L}_{\delta}:=(T_{\delta}-H)\partial_{x}.$
We may write the nonlinear term as
(2.1) $\mathcal{N}(u)$ $=-\beta u\mathcal{L}_{\delta}|u|^{2}+iu(\alpha+i\beta H)\partial_{x}|u|^{2}$
$=:\mathcal{N}_{\delta}(u)+\mathcal{N}_{\infty}(u).$
We have already defined the space $\mathcal{Z}^{k}$, the distance function $d^{k}(\cdot,\cdot)$ and the functional $\mathcal{E}^{k}(\cdot)$ on it, in (1.2), (1.3) and (1.4), respectively. We also define the smooth subset $\mathcal{Z}^{\infty}:=\bigcap_{k\geq 1}\mathcal{Z}^{k}$.
2.2. Properties of $T_{\delta}$
Having an unbounded symbol (1.1), the operator $T_{\delta}$ itself is not bounded on $L^{p}(\mathbb{R})$ or $H^{s}(\mathbb{R})$. However, since the Fourier symbol $r_{\delta}(\xi)=(\coth(\delta\xi)-\mathrm{sgn}(\xi))\xi$ of $\mathcal{L}_{\delta}$ satisfies $|r_{\delta}(\xi)|\lesssim\delta^{-1}$, by Plancherel’s theorem we see that $\mathcal{L}_{\delta}$ is bounded on $H^{s}(\mathbb{R})$ for any $s\in\mathbb{R}$ and satisfies
(2.2) $\|\mathcal{L}_{\delta}f\|_{H^{s}}\lesssim\delta^{-1}\|f\|_{H^{s}}.$
Combining (2.2) with the boundedness of $H$, we also see that $T_{\delta}\partial_{x}$ is bounded from $H^{s+1}(\mathbb{R})$ to $H^{s}(\mathbb{R})$ and satisfies
(2.3) $\|T_{\delta}\partial_{x}f\|_{H^{s}}\lesssim\|\partial_{x}f\|_{H^{s}}+\delta^{-1}\|f\|_{H^{s}}.$
For $f\in\mathcal{Z}^{k}$, we notice that $\partial_{x}|f|^{2}=\partial_{x}(|f|^{2}-\rho_{f}^{2})$ gives
(2.4) $\|T_{\delta}\partial_{x}|f|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{2}},$
$\|\mathcal{L}_{\delta}|f|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}.$
We also remark that the operator $T_{\delta}\partial_{x}$ is positive, symmetric, and satisfies $\overline{T_{\delta}\partial_{x}f}=T_{\delta}\partial_{x}\overline{f}$:
(2.5) $\left(T_{\delta}\partial_{x}f,f\right)_{L^{2}}\geq 0,\quad\left(T_{\delta}\partial_{x}f,g\right)_{L^{2}}=\left(f,T_{\delta}\partial_{x}g\right)_{L^{2}},\qquad\int_{\mathbb{R}}fT_{\delta}\partial_{x}g\,dx=\int_{\mathbb{R}}gT_{\delta}\partial_{x}f\,dx,$
which follows from the fact that the Fourier symbol $\coth(\delta\xi)\xi$ of $T_{\delta}\partial_{x}$ is non-negative and even. The same is true for the operator $H\partial_{x}$.
2.3. Properties of the space $\mathcal{Z}^{k}$ and control of Littlewood-Paley operators
We begin this subsection with a basic estimate in $\mathcal{Z}^{k}$, which will be used in Section 4:
(2.6) $\rho\leq\|f\|_{L^{\infty}}\lesssim\rho+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{1}{3}}\|\partial_{x}f\|_{L^{2}}^{\frac{1}{3}}\qquad(f\in\mathcal{Z}^{1}_{\rho},~\rho\geq 0).$
The first inequality follows from $|f|^{2}-\rho^{2}\in L^{2}(\mathbb{R})$. The second inequality is obtained by Gagliardo-Nirenberg and Young’s inequalities as follows:
$\|f\|_{L^{\infty}}^{2}$ $\leq\rho^{2}+\big\||f|^{2}-\rho^{2}\big\|_{L^{\infty}}\leq\rho^{2}+C\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{1}{2}}\|f\|_{L^{\infty}}^{\frac{1}{2}}\|\partial_{x}f\|_{L^{2}}^{\frac{1}{2}}$
$\leq\rho^{2}+\frac{1}{2}\|f\|_{L^{\infty}}^{2}+C\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{2}{3}}\|\partial_{x}f\|_{L^{2}}^{\frac{2}{3}}.$
Next, from Sobolev’s embedding $H^{1}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$ we see that for any integer $k\geq 1$, $\rho\geq 0$, $f\in\mathcal{Z}^{k}_{\rho}$ and $g\in H^{k}(\mathbb{R})$, it holds $f+g\in\mathcal{Z}^{k}_{\rho}$ and
(2.7) $\mathcal{E}^{k}(f+g)\lesssim\mathcal{E}^{k}(f)+\|g\|_{H^{k}}^{2}.$
In a similar manner (but using the Gagliardo-Nirenberg inequality instead of Sobolev’s embedding), for $f_{1},f_{2}\in\mathcal{Z}^{1}$ and $g_{1},g_{2}\in H^{1}(\mathbb{R})$, we can show that
(2.8) $d^{1}(f_{1}+g_{1},f_{2}+g_{2})$ $\lesssim\big(1+\|g_{1}\|_{L^{2}}\big)d^{1}(f_{1},f_{2})$
$\quad+\big(1+\|f_{2}\|_{L^{\infty}}+\|g_{1}\|_{L^{\infty}}+\|g_{2}\|_{L^{\infty}}\big)\|g_{1}-g_{2}\|_{L^{2}}+\|\partial_{x}(g_{1}-g_{2})\|_{L^{2}}.$
Below, we collect some estimates for frequency localized functions. They are easily proven by using Bernstein’s inequality. For any $\rho\geq 0$, $\ell\geq 0$ and $k,j\geq 1$, we can show that $P_{\leq\ell}:\mathcal{Z}^{k}_{\rho}\to\mathcal{Z}^{k+j}_{\rho}$ with the estimates
(2.9) $\begin{aligned} \mathcal{E}^{k}(P_{\leq\ell}f)&\leq\mathcal{E}^{k}(f)+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}\leq\mathcal{E}^{k}(f)+2\|f\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}\\ &\leq\mathcal{E}^{k}(f)+2^{1-\ell}\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{2}}\leq\mathcal{E}^{k}(f)+2^{1-\ell}\mathcal{E}^{1}(f)\leq 3\mathcal{E}^{k}(f),\end{aligned}$
$\mathcal{E}^{k+j}(P_{\leq\ell}f)\lesssim 2^{2j\ell}\mathcal{E}^{k}(f)+2^{-\ell}\mathcal{E}^{1}(f)\lesssim 2^{2j\ell}\mathcal{E}^{k}(f).$
We also see that $P_{\leq\ell}f\to f$ in $\mathcal{Z}^{k}$ as $\ell\to\infty$, thanks to the estimate
(2.10) $d^{k}(P_{\leq\ell}f,f)$ $\leq\|P_{\geq\ell+1}\partial_{x}f\|_{H^{k-1}}+\|P_{\geq\ell+1}f\|_{L^{\infty}}+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}$
$\lesssim\|P_{\geq\ell+1}\partial_{x}f\|_{H^{k-1}}+2^{-\frac{1}{2}\ell}\|P_{\geq\ell+1}\partial_{x}f\|_{L^{2}}+2^{-\ell}\|f\|_{L^{\infty}}\|P_{\geq\ell+1}\partial_{x}f\|_{L^{2}}.$
Note that if $f\in C([0,T];\mathcal{Z}^{k})$, then the above convergence is uniform in time on $[0,T]$. Using the identity
(2.11) $|Af|^{2}-|Ag|^{2}=|f|^{2}-|g|^{2}$ $-(1-A)(f-g)\cdot\overline{Af}-(1-A)g\cdot\overline{A(f-g)}$
$-(f-g)\cdot\overline{(1-A)f}-g\cdot\overline{(1-A)(f-g)}$
with $A=P_{\leq\ell}$, we have
(2.12) $\begin{aligned} d^{k}(P_{\leq\ell}f,P_{\leq\ell}g)&\leq d^{k}(f,g)\begin{aligned} &+\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\|P_{\leq\ell}f\|_{L^{\infty}}+\|P_{\geq\ell+1}g\|_{L^{2}}\|P_{\leq\ell}(f-g)\|_{L^{\infty}}\\ &+\|f-g\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}+\|g\|_{L^{\infty}}\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\end{aligned}\\ &\leq d^{k}(f,g)+2^{1-\ell}\Big(\big[\mathcal{E}^{1}(f)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}(g)\big]^{\frac{1}{2}}\Big)d^{1}(f,g).\end{aligned}$
Finally, for different $\ell,m\in\mathbb{Z}_{\geq 0}$ and $f,g\in\mathcal{Z}^{2}$, we use the triangle inequality and (2.10) with $k=1$ to see that
(2.13) $d^{1}(P_{\leq\ell}f,P_{\leq m}g)\lesssim d^{1}(f,g)+2^{-\ell}\Big(\big[\mathcal{E}^{2}(f)\big]^{\frac{1}{2}}+\mathcal{E}^{2}(f)\Big)+2^{-m}\Big(\big[\mathcal{E}^{2}(g)\big]^{\frac{1}{2}}+\mathcal{E}^{2}(g)\Big).$
2.4. Estimates for the free Schrödinger operator
The action of $U(t)$ in the Zhidkov-type spaces has been well-studied; see, e.g., [18, 11].
A standard Fourier analysis with the pointwise bound $|(e^{it\xi^{2}}-1)/\xi|\lesssim|t|^{1/2}$ reveals that
(2.14) $\|U(t)\phi-\phi\|_{L^{2}}\lesssim|t|^{\frac{1}{2}}\|\partial_{x}\phi\|_{L^{2}}.$
Using Gagliardo-Nirenberg inequality and (2.14), we see that
(2.15) $\|U(t)\phi-\phi\|_{L^{\infty}}$ $\lesssim|t|^{\frac{1}{4}}\|\partial_{x}\phi\|_{L^{2}},$
$\|U(t)\phi\|_{L^{\infty}}$ $\lesssim\|\phi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}\phi\|_{L^{2}},$
and by (2.14) and (2.15),
(2.16) $\big\||U(t)\phi|^{2}-|\phi|^{2}\big\|_{L^{2}}$ $\lesssim\Big(\|\phi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}\phi\|_{L^{2}}\Big)|t|^{\frac{1}{2}}\|\partial_{x}\phi\|_{L^{2}},$
$\big\||U(t)\phi|^{2}-\rho^{2}\big\|_{L^{2}}$ $\lesssim\big\||\phi|^{2}-\rho^{2}\big\|_{L^{2}}+\Big(\|\phi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}\phi\|_{L^{2}}\Big)|t|^{\frac{1}{2}}\|\partial_{x}\phi\|_{L^{2}}.$
These estimates with the unitarity of $U(t)$ in $H^{s}(\mathbb{R})$ imply that
(2.17) $U(t)\phi\in C(\mathbb{R};\mathcal{Z}^{k}_{\rho}),\quad\mathcal{E}^{k}(U(t)\phi)\lesssim(1+|t|^{\frac{3}{4}})\mathcal{E}^{k}(\phi)\qquad(\rho\geq 0,~\phi\in\mathcal{Z}^{k}_{\rho}).$
Finally, we see the continuity of the mapping $\phi\mapsto U(t)\phi$ by showing the difference estimate
(2.18) $d^{k}(U(t)\phi,U(t)\psi)\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}(\phi)+\mathcal{E}^{1}(\psi)\Big)^{\frac{1}{2}}d^{k}(\phi,\psi)$
for any $k\geq 1$ and $\phi,\psi\in\mathcal{Z}^{k}$. To prove (2.18), it is clearly sufficient to estimate the non-Sobolev part of $d^{1}$. From (2.15), we have
$\|U(t)\phi-U(t)\psi\|_{L^{\infty}}\lesssim\|\phi-\psi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}(\phi-\psi)\|_{L^{2}}\lesssim(1+|t|^{\frac{1}{4}})d^{1}(\phi,\psi).$
Using the decomposition (2.11) with $A=U(t)$ and (2.14)–(2.15), we have
$\big\|(|U(t)\phi|^{2}-\rho_{\phi}^{2})-(|U(t)\psi|^{2}-\rho_{\psi}^{2})\big\|_{L^{2}}$
$\lesssim\big\|(|\phi|^{2}-\rho_{\phi}^{2})-(|\psi|^{2}-\rho_{\psi}^{2})\big\|_{L^{2}}\begin{aligned} &+\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\|U(t)\phi\|_{L^{\infty}}+\|U(t)\psi-\psi\|_{L^{2}}\|U(t)(\phi-\psi)\|_{L^{\infty}}\\ &+\|\phi-\psi\|_{L^{\infty}}\|U(t)\phi-\phi\|_{L^{2}}+\|\psi\|_{L^{\infty}}\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\end{aligned}$
$\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}(\phi)+\mathcal{E}^{1}(\psi)\Big)^{\frac{1}{2}}d^{1}(\phi,\psi),$
which shows (2.18).
2.5. Nonlinear estimates and an integration by parts lemma
We give some nonlinear estimates in the following lemma.
Lemma 2.1.
  1. $\mathrm{(i)}$
    For $f\in\mathcal{Z}^{1}$,
    $\|\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim\delta^{-1}\big[\mathcal{E}^{1}(f)\big]^{\frac{3}{2}}.$
    For any integer $k\geq 1$ and $f\in\mathcal{Z}^{k}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim_{k}\delta^{-1}\mathcal{E}^{1}(f)\|\partial_{x}f\|_{H^{k-1}}.$
  2. $\mathrm{(ii)}$
    For any integer $k\geq 0$ and $f\in\mathcal{Z}^{k+1}_{*}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\infty}(f)\|_{L^{2}}\lesssim_{k}\mathcal{E}^{1}(f)\|\partial_{x}f\|_{H^{k}}.$
  3. $\mathrm{(iii)}$
    For $f_{1},f_{2}\in\mathcal{Z}^{1}$,
    $\delta\;\|\mathcal{N}_{\delta}(f_{1})-\mathcal{N}_{\delta}(f_{2})\|_{H^{1}}+\|\mathcal{N}_{\infty}(f_{1})-\mathcal{N}_{\infty}(f_{2})\|_{L^{2}}\lesssim\Big(1+\mathcal{E}^{1}(f_{1})+\mathcal{E}^{1}(f_{2})\Big)d^{1}(f_{1},f_{2}).$
Proof.
We use the boundedness of $\mathcal{L}_{\delta}$ ((2.2), (2.4)) and $H$ on $L^{2}(\mathbb{R})$ for estimates of $\mathcal{N}_{\delta}$ and $\mathcal{N}_{\infty}$, respectively. In the case $k\geq 2$, we use appropriate interpolation inequalities to bound $\|\partial_{x}^{j}f\|_{L^{2}}$ and $\|\partial_{x}^{j}f\|_{L^{\infty}}$ for $1\leq j\leq k-1$ in terms of $\|\partial_{x}f\|_{L^{2}}$ and $\|\partial_{x}^{k}f\|_{L^{2}}$. A detailed verification is left to the reader. ∎
We conclude this section with the following lemma, which describes typical integration-by-parts procedure in the energy method.
Lemma 2.2.
Let $f\in W^{1,\infty}(\mathbb{R})$, and $g\in H^{1}(\mathbb{R})$. The following estimate holds:
$\big|\operatorname{Re}\big(f(\overline{f}\partial_{x}g),g\big)_{L^{2}}\big|+\big|\operatorname{Re}\big(f(f\partial_{x}\overline{g}),g\big)_{L^{2}}\big|+\big|\operatorname{Re}\big(f(iH)(\overline{f}\partial_{x}g),g\big)_{L^{2}}\big|\lesssim\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{\infty}}\|g\|_{L^{2}}^{2}.$
Remark 2.1.
The right-hand side does not include $\partial_{x}g$. We do not have a similar estimate for the term $\operatorname{Re}\big(f(iH)(f\partial_{x}\overline{g}),g\big)_{L^{2}}$, which suggests that the standard energy method is not sufficient to control this type of interactions. We will use the modified energy method to handle this case.
Proof of Lemma 2.2.
The first two terms are easily treated by integration by parts (one can estimate the second term without taking the real part). For the term involving $H$, we rewrite it as
$\operatorname{Re}\big(f(iH)(\overline{f}\partial_{x}g),g\big)_{L^{2}}=\operatorname{Re}\big(iH\partial_{x}(\overline{f}g),\overline{f}g\big)_{L^{2}}-\operatorname{Re}\big(f(iH)(g\partial_{x}\overline{f}),g\big)_{L^{2}}.$
Observe that the first term on the right-hand side vanishes by the positivity of $H\partial_{x}$, then we obtain the desired bound. ∎
3. Local well-posedness
In this section, we prove our main local well-posedness result, Theorem 1.1. Note that Corollary 1.2 can be proven by the same argument but omitting the nonlinear term $\mathcal{N}_{\delta}(\cdot)$ in the calculations. At the end, we will give a proof of Corollary 1.3. Throughout this section, we fix $\alpha,\beta\in\mathbb{R}$ and do not indicate the dependence of constants on these parameters.
We employ a compactness argument. For $\ell\in\mathbb{Z}_{\geq 0}$, consider the mollified equation for (gINLS) with the mollification parameter $\ell$:
(3.1) $i\partial_{t}u-\partial_{x}^{2}u=P_{\leq\ell}\mathcal{N}(P_{\leq\ell}u).$
The associated integral equation is given by
(3.2) $u=U(\cdot)u(0)+\Phi^{(\ell)}[u],\qquad\Phi^{(\ell)}[u](t):=-i\int_{0}^{t}U(t-s)P_{\leq\ell}\mathcal{N}(P_{\leq\ell}u(s))\,ds.$
The proof proceeds in four main steps. First, in Proposition 3.1, we establish the existence of a unique local solution $u^{(\ell)}\in C([0,T_{\ell});\mathcal{Z}^{\infty})$ to the mollified problem (3.1) with mollified initial condition $u^{(\ell)}(0)=P_{\leq\ell}\phi$. Second, in Proposition 3.2, we derive uniform-in-$\ell$ existence time $T_{0}>0$ and $\mathcal{Z}^{2}$ a priori bound for the sequence $\{u^{(\ell)}\}_{\ell\geq 0}$, valid for all $t\in[0,T_{0}]$. Third, in Proposition 3.3, we prove that $\{u^{(\ell)}\}_{\ell\geq 0}$ is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1})$ for some possibly shorter time $T_{*}\leq T_{0}$, which gives us a solution of (gINLS) on $0,T_{*}]$ by taking the limit $\ell\to\infty$. Finally, we prove uniqueness and continuity of the flow map in $\mathcal{Z}^{2}$, completing the proof of Theorem 1.1.
Proposition 3.1.
Let $\phi\in\mathcal{Z}^{1}$. For each $\ell\in\mathbb{Z}_{\geq 0}$, the Cauchy problem (3.1) with $u(0)=P_{\leq\ell}\phi$ admits a unique maximal solution
$u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}^{\infty}),$
where $T_{\ell}(\phi)$ is the maximal existence time of $u^{(\ell)}$ for which the following blow-up criterion holds: if $T_{\ell}(\phi)<\infty$, then
$\lim_{t\uparrow T_{\ell}(\phi)}\mathcal{E}^{1}(u^{(\ell)}(t))=\infty.$
Proof of Proposition 3.1.
It is sufficient to construct a unique local solution $u^{(\ell)}\in C([0,T];\mathcal{Z}^{1})$ to the initial value problem of (3.1) with initial data $\phi$ satisfying $\mathrm{supp}\,(\widehat{\phi})\subset\{|\xi|\leq 2^{\ell+1}\}$ and $\mathcal{E}^{1}(\phi)\leq M_{0}^{2}$, for some $T>0$ depending on $M_{0}$ and $\ell$. Indeed, the solution $u^{(\ell)}$ then satisfies $\widehat{u^{(\ell)}}(\xi,t)=0$ outside $\{|\xi|\leq 2^{\ell+1}\}$ for all $t\geq 0$, from which the regularity of $u^{(\ell)}$ follows.
Recall from (2.7) and (2.17) that $\|\mathcal{E}^{1}(U(t)\phi+\Phi^{(\ell)}(u))\|_{L^{\infty}(0,1)}\leq C_{*}\mathcal{E}^{1}(\phi)+C\|\Phi^{(\ell)}(u)\|_{L^{\infty}(0,1;H^{1})}^{2}$ for some $C_{*},C>0$. For $0<T\leq 1$, define the complete metric space $(X_{T},d_{T})$ by
$X_{T}:=\big\{u\in C([0,T];\mathcal{Z}^{1}):\|\mathcal{E}^{1}(u)\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}\big\},\qquad d_{T}(u,v):=\|d^{1}(u,v)\|_{L^{\infty}_{T}},$
and we shall show that the mapping $\mathcal{S}^{(\ell)}_{\phi}:u\mapsto U(\cdot)\phi+\Phi^{(\ell)}[u]$ is a contraction on $X_{T}$.
By Bernstein’s inequality, Lemma 2.1, and (2.9), for $u\in X_{T}$ we obtain
(3.3) $\|\Phi^{(\ell)}[u]\|_{L^{\infty}_{T}H^{1}}\lesssim T\big(\|\mathcal{N}_{\delta}(P_{\leq\ell}u)\|_{L^{\infty}_{T}H^{1}}+2^{\ell}\|\mathcal{N}_{\infty}(P_{\leq\ell}u)\|_{L^{\infty}_{T}L^{2}}\big)\lesssim T(\delta^{-1}+2^{\ell})M_{0}^{3},$
and similarly but using (2.12) instead of (2.9), for $u,v\in X_{T}$
(3.4) $\|\Phi^{(\ell)}[u]-\Phi^{(\ell)}[v]\|_{L^{\infty}_{T}H^{1}}\lesssim T(\delta^{-1}+2^{\ell})(1+M_{0}^{3})d_{T}(u,v).$
By (2.7) and (3.3), we have $\|\mathcal{E}^{1}(\mathcal{S}^{(\ell)}_{\phi}[u])\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}$ if $T>0$ is chosen as $T(\delta^{-1}+2^{\ell})M_{0}^{2}\ll 1$. Moreover, using (2.8), (2.18) and (2.14), we can show that the mapping $t\mapsto\mathcal{S}^{(\ell)}_{\phi}[u](t)$ is continuous in $\mathcal{Z}^{1}$, and hence $\mathcal{S}^{(\ell)}_{\phi}$ maps $X_{T}$ into itself. Furthermore, (2.8) with (2.15), (3.3) and (3.4) imply that
$d_{T}(\mathcal{S}^{(\ell)}_{\phi}[u],\mathcal{S}^{(\ell)}_{\phi}[v])$ $\lesssim\big(1+\|U(t)\phi\|_{L^{\infty}_{T}L^{\infty}}+\|\Phi^{(\ell)}[u]\|_{L^{\infty}_{T}L^{\infty}}+\|\Phi^{(\ell)}[v]\|_{L^{\infty}_{T}L^{\infty}}\big)\|\Phi^{(\ell)}[u]-\Phi^{(\ell)}[v]\|_{L^{\infty}_{T}H^{1}}$
$\lesssim\big(1+M_{0}+T(\delta^{-1}+2^{\ell})M_{0}^{3}\big)T(\delta^{-1}+2^{\ell})(1+M_{0}^{3})d_{T}(u,v).$
Therefore, $\mathcal{S}^{(\ell)}_{\phi}$ is a contraction on $X_{T}$ if $T$ is chosen as $T(\delta^{-1}+2^{\ell})(1+M_{0}^{4})\ll 1$, and the contraction mapping theorem shows that there exists a unique solution $u^{(\ell)}\in X_{T}$ to the initial value problem. The uniqueness in the whole space $C([0,T];\mathcal{Z}^{1})$ and the blow-up criterion can be shown by a standard argument; we have thus proven the proposition. ∎
Proposition 3.2 (Uniform $\mathcal{Z}^{2}$ estimate).
Let $M_{0}>0$, and $\phi\in\mathcal{Z}^{2}$ be such that $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$. For $\ell\geq 0$, let $u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}^{\infty})$ be the unique forward maximal solution to (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ given in Proposition 3.1. Then, there exists $\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}\lesssim T_{0}\leq 1$ such that for all $\ell$, we have $T_{\ell}(\phi)>T_{0}$ and
(3.5) $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim\mathcal{E}^{2}(\phi).$
Moreover, for any integer $k\geq 3$ we have
(3.6) $\|\mathcal{E}^{k}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim_{k}\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}(\phi).$
Proof.
To establish the uniform bound (3.5), the principal difficulty arises from a derivative loss in the highest-order energy estimate. To overcome this, we employ a modified energy method: we design a frequency-localized correction term that neutralizes the problematic nonlinear interactions in the nonlinearity of (3.1).
Let $\ell\geq 0$, and take arbitrary $T\in(0,1]$ with $T<T_{\ell}(\phi)$. In the following, we put
$\zeta_{\ell}:=P_{\leq\ell}u^{(\ell)}.$
We begin by controlling the lower-order part $\mathcal{E}^{1}(u^{(\ell)})$ using the integral equation (3.2) and (2.7) with $k=1$,
(3.7) $\|\mathcal{E}^{1}(u^{(\ell)})\|_{L^{\infty}_{T}}$ $\lesssim\|\mathcal{E}^{1}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}H^{1}}^{2}$
$\leq\|\mathcal{E}^{1}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\Big(\int_{0}^{T}\|\mathcal{N}(\zeta_{\ell}(s))\|_{H^{1}}\,ds\Big)^{2}$
$\lesssim\mathcal{E}^{1}(\phi)+T^{2}(1+\delta^{-1})^{2}\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3},$
where we have used (2.17), (2.9) and Lemma 2.1 to derive the last inequality.
Now, we move on to the energy estimate for the highest-order component of $\mathcal{E}^{2}(u^{(\ell)})$. Note that the following formal computations are justified since $u^{(\ell)}$ is a smooth solution. Using the equation (3.1), we find
$\frac{d}{dt}\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}$ $=2\operatorname{Re}\big(\partial_{x}^{2}\partial_{t}u^{(\ell)}(t),\partial_{x}^{2}u^{(\ell)}(t)\big)_{L^{2}}=2\operatorname{Re}\big({-}i\partial_{x}^{2}\mathcal{N}(\zeta_{\ell}(t)),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}$
$=2\operatorname{Re}\big({-}i\partial_{x}^{2}\mathcal{N}_{\delta}(\zeta_{\ell}(t)),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}+2\operatorname{Re}\big(\partial_{x}^{2}\big(\zeta_{\ell}(t)(\alpha+i\beta H)\partial_{x}|\zeta_{\ell}(t)|^{2}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}.$
The first term is harmless and can be estimated directly by using Lemma 2.1,
$\Big|\big({-}i\partial_{x}^{2}\mathcal{N}_{\delta}(\zeta_{\ell}(t)),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|\lesssim\delta^{-1}\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
For the second term, after distributing the differentials we easily estimate in the case where no function has three differentials; that is,
$\Big|2\operatorname{Re}\big(\partial_{x}^{2}\big(\zeta_{\ell}(t)(\alpha+i\beta H)\partial_{x}|\zeta_{\ell}(t)|^{2}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}-2\operatorname{Re}\big(\zeta_{\ell}(t)(\alpha+i\beta H)\big(\partial_{x}^{3}\zeta_{\ell}(t)\,\overline{\zeta_{\ell}(t)}+\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|$
$\lesssim\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
Even for the remaining case, a standard integration by parts argument with Lemma 2.2 can be applied except for one term; that is,
$\Big|2\operatorname{Re}\big(\zeta_{\ell}(t)(\alpha+i\beta H)\big(\partial_{x}^{3}\zeta_{\ell}(t)\,\overline{\zeta_{\ell}(t)}+\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}-2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|$
$\lesssim\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
To summarize, we have obtained
(3.8) $\frac{d}{dt}\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
The first term on the right-hand side of (3.8) is problematic, as it contains a derivative loss that cannot be handled by the standard energy method. We introduce a time-dependent corrector to compensate with this term. For a frequency threshold $n_{0}\in\mathbb{Z}_{\geq 0}$ to be chosen sufficiently large later, define
(3.9) $\mathcal{I}_{2}(t):=\beta\operatorname{Re}\big(\zeta_{\ell}(t)H\partial_{x}\big(\zeta_{\ell}(t)\partial_{x}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}.$
This corrector is designed so that its time derivative cancels the problematic term. By Bernstein’s inequality, we obtain the bound
(3.10) $|\mathcal{I}_{2}(t)|$ $\lesssim\|\zeta_{\ell}(t)\|_{L^{\infty}}\|\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\|_{L^{2}}\|\zeta_{\ell}(t)\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\|_{H^{1}}$
$\lesssim 2^{-n_{0}}\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}.$
Differentiating $\mathcal{I}_{2}(t)$ and using the equation (3.1) and the symmetry property (2.5) for $H\partial_{x}$, we see that
$\frac{d}{dt}\mathcal{I}_{2}(t)$ $=2\beta\operatorname{Re}\big(H\partial_{x}\big(\partial_{t}\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}$
$\quad+2\beta\operatorname{Re}\big(H\partial_{x}\big(\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}\partial_{t}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}$
$=2\beta\operatorname{Re}\big(iH\partial_{x}\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}+\mathcal{R}(t),$
where the remainder term $\mathcal{R}(t)$ collects all terms having at most two differentials on each function,
$\mathcal{R}(t)$ $=-2\beta\operatorname{Re}\big({-}iH\big(\partial_{x}^{2}\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\partial_{x}\big(\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)\big)_{L^{2}}$
$\quad+2\beta\operatorname{Re}\big({-}iH\partial_{x}\big(P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))\cdot\partial_{x}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}$
$\quad-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))}\big),\partial_{x}\big(\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)\big)_{L^{2}}.$
Using Lemma 2.1, we see that
$|\mathcal{R}(t)|\lesssim\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
Moreover, the main term in $\frac{d}{dt}\mathcal{I}_{2}(t)$ compensates with the first term in (3.8), that is,
$\Big|2\beta\operatorname{Re}\big(iH\partial_{x}\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}+2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|$
$\leq\Big|2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\partial_{x}\overline{\zeta_{\ell}(t)}\cdot\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}\Big|$
$\quad+\Big|2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{P_{\geq n_{0}}\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}P_{\geq n_{0}}\zeta_{\ell}(t)\big)_{L^{2}}-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|$
$\lesssim(1+2^{n_{0}})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2},$
where we have applied integration by parts several times and used Bernstein’s inequality as
$\|\partial_{x}^{3}P_{\leq n_{0}-1}\zeta_{\ell}\|_{L^{2}}\lesssim 2^{n_{0}}\|\partial_{x}^{2}u^{(\ell)}\|_{L^{2}}.$
To summarize, we have obtained
(3.11) $\frac{d}{dt}\mathcal{I}_{2}(t)$ $\leq-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}$
$\quad+C(1+2^{n_{0}})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
Combining (3.8) and (3.11), we obtain the modified energy inequality,
$\frac{d}{dt}\big(\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{2}(t)\big)$ $\leq C(1+\delta^{-1}+2^{n_{0}})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
Then, integrating it from 0 to $t$, and using the bound (3.10) of $\mathcal{I}_{2}(t)$, we see that
(3.12) $\|\partial_{x}^{2}u^{(\ell)}\|_{L^{\infty}_{T}L^{2}}^{2}\leq\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+C\Big\{\big(2^{-n_{0}}+T(1+\delta^{-1}+2^{n_{0}})\big)\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}+T(1+\delta^{-1})\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3}\Big\}.$
Putting (3.7) and (3.12) together, we obtain
(3.13) $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}(\phi)+\Big\{2^{-n_{0}}\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}+T(1+\delta^{-1}+2^{n_{0}})\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}$
$+\big[T(1+\delta^{-1})+T^{2}(1+\delta^{-1})^{2}\big]\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}\Big\}$ $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}.$
We now choose $n_{0}$ as $2^{n_{0}}=C(1+M_{0}^{2})$ with a large constant $C>0$ so that $2^{-n_{0}}\mathcal{E}^{2}(\phi)\leq 2^{-n_{0}}M_{0}^{2}\ll 1$, and choose $T_{0}$ sufficiently small so that $T_{0}(1+\delta^{-1})\big((1+2^{n_{0}})M_{0}^{2}+M_{0}^{4}\big)\ll 1$. Then, (3.13) and a standard continuity argument show that
$\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}(\phi),\qquad 0<T<\min\{T_{0},T_{\ell}(\phi)\}.$
By the blow-up criterion in Proposition 3.1, we see that $T_{\ell}(\phi)>T_{0}$ and that (3.5) holds.
The a priori bound in higher regularity (3.6) can be derived by a similar modified energy argument; we only outline the proof. It suffices to estimate the highest component $\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}$. Arguing similarly to the proof of (3.5), by Lemmas 2.1, 2.2 and interpolation inequalities, we deduce that
(3.14) $\frac{d}{dt}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}+C_{k}(1+\delta^{-1})\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}.$
Define the corrector as
$\mathcal{I}_{k}(t):=\beta\operatorname{Re}\big(\zeta_{\ell}(t)H\partial_{x}\big(\zeta_{\ell}(t)\partial_{x}^{k-1}\overline{P_{\geq n_{1}}\zeta_{\ell}(t)}\big),\partial_{x}^{k-1}P_{\geq n_{1}}\zeta_{\ell}(t)\big)_{L^{2}}$
for $n_{1}\gg 1$ to be chosen later. Then we have
$|\mathcal{I}_{k}(t)|\lesssim_{k}2^{-n_{1}}\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},$
$\begin{aligned} \frac{d}{dt}\mathcal{I}_{k}(t)&=-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}\\ &\quad+C_{k}\Big\{(1+2^{n_{1}})\mathcal{E}^{2}(u^{(\ell)}(t))+(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}\Big\}\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}\end{aligned}$
in analogy with (3.10) and (3.11).
Now, we invoke the $\mathcal{Z}^{2}$ bound (3.5) and choose $n_{1}$ as $2^{n_{1}}=C(1+M_{0}^{2})$ with $C\gg_{k}1$. Then, we have
$|\mathcal{I}_{k}(t)|\leq\frac{1}{2}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},\qquad t\in[0,T_{0}],$
and the modified energy inequality
$\frac{d}{dt}\big(\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)\big)\lesssim_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}(\phi)+\big[\mathcal{E}^{2}(\phi)\big]^{2}\Big)\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2},\qquad t\in[0,T_{0}].$
Noticing that $\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}\lesssim\mathcal{E}^{1}(u^{(\ell)}(t))+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\lesssim\mathcal{E}^{2}(\phi)+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}$, we see that the function
$X_{k}(t):=\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)+\mathcal{E}^{2}(\phi)$
satisfies the differential inequality
$\frac{d}{dt}X_{k}(t)\leq C_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}(\phi)+\big[\mathcal{E}^{2}(\phi)\big]^{2}\Big)X_{k}(t),\qquad t\in[0,T_{0}].$
Using Gronwall’s inequality and recalling $T_{0}(1+\delta^{-1})((1+2^{n_{0}})M_{0}^{2}+M_{0}^{4})\ll 1$, we obtain
$\|X_{k}\|_{L^{\infty}_{T_{0}}}\leq X_{k}(0)e^{C_{k}T_{0}(1+\delta^{-1})\big(\mathcal{E}^{2}(\phi)+[\mathcal{E}^{2}(\phi)]^{2}\big)}\leq Ce^{C_{k}}\Big(\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}(\phi)\Big).$
Since $\mathcal{E}^{k}(u^{(\ell)}(t))\lesssim X_{k}(t)$ for $t\in[0,T_{0}]$ by (3.5), this yields (3.6). ∎
Proposition 3.3 ($\mathcal{Z}^{1}$ difference estimate).
Let $\delta>0$, $M_{0}>0$, and let $\phi,\psi\in\mathcal{Z}^{2}$ satisfy $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$ and $\mathcal{E}^{2}(\psi)\leq M_{0}^{2}$. For $\ell,m\geq 0$, let $u^{(\ell)},v^{(m)}\in C([0,T_{0}];\mathcal{Z}^{\infty})$ be the solutions to (3.1) with the mollification parameters $\ell,m$ and the initial data $P_{\leq\ell}\phi$ and $P_{\leq m}\psi$, respectively, where $T_{0}\in(0,1]$ is the constant given in Proposition 3.2. Then, there exists $T_{*}\in(0,T_{0}]$ with $T_{*}\gtrsim\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ such that we have
(3.15) $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
In particular, if $\phi=\psi$, then
(3.16) $\|d^{1}(u^{(\ell)},u^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
Proof.
In the following, we write
$\zeta_{\ell}:=P_{\leq\ell}u^{(\ell)},\quad\eta_{m}:=P_{\leq m}v^{(m)},\quad\omega_{\ell,m}:=\zeta_{\ell}-\eta_{m},\quad\mu_{\ell,m}:=\frac{\zeta_{\ell}+\eta_{m}}{2}.$
Take an arbitrary $T\in(0,T_{0}]$ with $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$. We see from Proposition 3.2 and (2.9) that
(3.17) $\|\mathcal{E}^{2}(\zeta_{\ell})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}(\eta_{m})\|_{L^{\infty}_{T}}\lesssim\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}(v^{(m)})\|_{L^{\infty}_{T}}\lesssim M_{0}^{2}.$
These bounds and Lemma 2.1 imply that
$\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}H^{1}}+\|\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}H^{1}}\lesssim T(1+\delta^{-1})M_{0}^{3}\lesssim 1.$
Moreover, we have the following difference estimates: by (2.18) and (2.13),
$\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}$ $\lesssim(1+M_{0})d^{1}(P_{\leq\ell}\phi,P_{\leq m}\psi)$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3}),$
and by Bernstein’s inequality, Lemma 2.1, and (2.13),
$\|\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{2}}$
$\lesssim T\Big(2^{-\ell}\|\mathcal{N}(\zeta_{\ell})\|_{L^{\infty}_{T}H^{1}}+2^{-m}\|\mathcal{N}(\eta_{m})\|_{L^{\infty}_{T}H^{1}}+\|\mathcal{N}(\zeta_{\ell})-\mathcal{N}(\eta_{m})\|_{L^{\infty}_{T}L^{2}}\Big)$
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{3})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}$
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{4})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}.$
Combining (2.8) and the above estimates, we obtain
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}L^{2}})\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}])\|_{L^{\infty}_{T}L^{2}}$
$\quad+\big(1+\|U(t)P_{\leq m}\psi\|_{L^{\infty}_{T}L^{\infty}}+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}L^{\infty}}+\|\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{\infty}}\big)\|\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{2}}$
$\lesssim\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}+(1+M_{0})\|\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{2}}$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}$
$\quad+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{5})+T(1+\delta^{-1})(1+M_{0}^{3})\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}},$
where at the second inequality we have used the equation as $\Phi^{(\ell)}[u^{(\ell)}]=u^{(\ell)}-U(t)P_{\leq\ell}\phi$. Imposing further smallness $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$ if necessary, we can absorb the last term into the left-hand side and obtain
(3.18) $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}.$
It remains to estimate the last term on the right-hand side of (3.18). Using the equation (3.1), we see that
$\frac{d}{dt}\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}$ $=2\operatorname{Re}\big({-}i\partial_{x}\big(P_{\leq\ell}\mathcal{N}(\zeta_{\ell}(t))-P_{\leq m}\mathcal{N}(\eta_{m}(t))\big),\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big)_{L^{2}}.$
We first rearrange the frequency projections by using Bernstein’s inequality, integration by parts, Lemma 2.1 and (3.17) as
$\frac{d}{dt}\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}=2\operatorname{Re}\big({-}i\partial_{x}\big(\mathcal{N}(\zeta_{\ell}(t))-\mathcal{N}(\eta_{m}(t))\big),\partial_{x}\big(\zeta_{\ell}(t)-\eta_{m}(t)\big)\big)_{L^{2}}+\mathcal{R}_{1}(t),$
with
$|\mathcal{R}_{1}(t)|\lesssim(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{4}.$
Next, we expand the difference of the nonlinearity as
$\mathcal{N}(\zeta_{\ell})-\mathcal{N}(\eta_{m})$ $=\mathcal{N}_{\delta}(\zeta_{\ell})-\mathcal{N}_{\delta}(\eta_{m})$
$\quad+\frac{i}{2}\omega_{\ell,m}(\alpha+i\beta H)\partial_{x}(|\zeta_{\ell}|^{2}+|\eta_{m}|^{2})$
$\quad+i\mu_{\ell,m}(\alpha+i\beta H)\partial_{x}(\overline{\mu_{\ell,m}}\omega_{\ell,m}+\mu_{\ell,m}\overline{\omega_{\ell,m}}),$
and hence
$2\operatorname{Re}\big({-}i\partial_{x}\big(\mathcal{N}(\zeta_{\ell}(t))-\mathcal{N}(\eta_{m}(t))\big),\partial_{x}\big(\zeta_{\ell}(t)-\eta_{m}(t)\big)\big)_{L^{2}}$
$=2\operatorname{Re}\big({-}i\partial_{x}\big(\mathcal{N}_{\delta}(\zeta_{\ell}(t))-\mathcal{N}_{\delta}(\eta_{m}(t))\big),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}$
$\quad+\operatorname{Re}\big(\partial_{x}\big(\omega_{\ell,m}(t)(\alpha+i\beta H)\partial_{x}(|\zeta_{\ell}(t)|^{2}+|\eta_{m}(t)|^{2})\big),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}$
$\quad+2\operatorname{Re}\big(\partial_{x}\big(\mu_{\ell,m}(t)(\alpha+i\beta H)\partial_{x}(\overline{\mu_{\ell,m}(t)}\omega_{\ell,m}(t)+\mu_{\ell,m}(t)\overline{\omega_{\ell,m}(t)})\big),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}.$
The contribution of the $\mathcal{N}_{\delta}$ part is easily estimated with Lemma 2.1 by
$\|\partial_{x}(\mathcal{N}_{\delta}(\zeta_{\ell})-\mathcal{N}_{\delta}(\eta_{m}))\|_{L^{\infty}_{T}L^{2}}\|\partial_{x}\omega_{\ell,m}\|_{L^{\infty}_{T}L^{2}}\lesssim\delta^{-1}(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
For the remaining terms, since $\omega_{\ell,m}$ corresponds to the difference $u^{(\ell)}-v^{(m)}$, by integration by parts, Lemma 2.2 and (3.17), we see that the only problematic interaction is
$2\beta\operatorname{Re}\big(\mu_{\ell,m}(t)(iH)(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{\omega_{\ell,m}(t)})),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}.$
Indeed, the others can be bounded by $M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}$. So far, we have proven that
(3.19) $\frac{d}{dt}\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}$ $\leq 2\beta\operatorname{Re}\big(\mu_{\ell,m}(t)(iH)(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{\omega_{\ell,m}(t)}),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}$
$\quad+C(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{4}+C(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
A corrector to cancel the problematic interaction in this case is given by
$\mathcal{J}(t):=\beta\operatorname{Re}\big(\mu_{\ell,m}(t)H\partial_{x}\big(\mu_{\ell,m}(t)\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),P_{\geq n_{2}}\omega_{\ell,m}(t)\big)_{L^{2}}$
with a large integer $n_{2}$ to be chosen later. Similar to the proof of Proposition 3.2, we see that
(3.20) $\frac{d}{dt}\mathcal{J}(t)$ $=-2\beta\operatorname{Re}\big(iH\big(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),\partial_{x}\big(\overline{\mu_{\ell,m}(t)}P_{\geq n_{2}}\omega_{\ell,m}(t)\big)\big)_{L^{2}}+\mathcal{R}_{2}(t)$
$=-2\beta\operatorname{Re}\big(iH\big(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{\omega_{\ell,m}(t)}\big),\overline{\mu_{\ell,m}(t)}\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}+\mathcal{R}_{2}(t)+\mathcal{R}_{3}(t),$
where
$\mathcal{R}_{2}(t)$ $:=-2\beta\operatorname{Re}\big({-}iH\big(\partial_{x}^{2}\mu_{\ell,m}(t)\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),\partial_{x}\big(\overline{\mu_{\ell,m}(t)}P_{\geq n_{2}}\omega_{\ell,m}(t)\big)\big)_{L^{2}}$
$\;\quad+\beta\operatorname{Re}\big({-}iH\partial_{x}\big((P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))+P_{\leq m}^{2}\mathcal{N}(\eta_{m}(t)))\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),\overline{\mu_{\ell,m}(t)}P_{\geq n_{2}}\omega_{\ell,m}(t)\big)_{L^{2}}$
$\;\quad-2\beta\operatorname{Re}\big(iH\big(\mu_{\ell,m}(t)\overline{P_{\geq n_{2}}(P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))-P_{\leq m}^{2}\mathcal{N}(\eta_{m}(t))}\big),\partial_{x}\big(\overline{\mu_{\ell,m}(t)}P_{\geq n_{2}}\omega_{\ell,m}(t)\big)\big)_{L^{2}}$
and
$\mathcal{R}_{3}(t)$ $:=2\beta\operatorname{Re}\big(iH\big(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{P_{\leq n_{2}-1}\omega_{\ell,m}(t)}\big),\partial_{x}\big(\overline{\mu_{\ell,m}(t)}P_{\geq n_{2}}\omega_{\ell,m}(t)\big)\big)_{L^{2}}$
$\;\quad+2\beta\operatorname{Re}\big(iH\big(\partial_{x}\mu_{\ell,m}(t)\cdot\partial_{x}\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),\partial_{x}\overline{\mu_{\ell,m}(t)}\cdot P_{\geq n_{2}}\omega_{\ell,m}(t)-\overline{\mu_{\ell,m}(t)}\partial_{x}P_{\leq n_{2}-1}\omega_{\ell,m}(t)\big)_{L^{2}}$
$\;\quad+2\beta\operatorname{Re}\big(iH\big(\mu_{\ell,m}(t)\partial_{x}\overline{P_{\geq n_{2}}\omega_{\ell,m}(t)}\big),\partial_{x}\big(\partial_{x}\overline{\mu_{\ell,m}(t)}\cdot P_{\geq n_{2}}\omega_{\ell,m}(t)-\overline{\mu_{\ell,m}(t)}\partial_{x}P_{\leq n_{2}-1}\omega_{\ell,m}(t)\big)\big)_{L^{2}}.$
Invoking the estimates obtained so far, we have
(3.21) $|\mathcal{R}_{2}(t)|$ $\lesssim(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{6}+(1+\delta^{-1})(M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
$|\mathcal{R}_{3}(t)|$ $\lesssim(1+2^{n_{2}})M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
Putting (3.19)–(3.21) together, we obtain the estimate
(3.22) $\frac{d}{dt}\Big(\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}+\mathcal{J}(t)\Big)$ $\lesssim(2^{-\ell}+2^{-m})(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
Integrating (3.22) from 0 to $t$, together with the estimate
$\|\mathcal{J}(t)\|_{L^{\infty}_{T}}\lesssim 2^{-n_{2}}M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
yields
(3.23) $\big\|\partial_{x}\big(u^{(\ell)}-v^{(m)}\big)\big\|_{L^{\infty}_{T}L^{2}}^{2}$ $\lesssim\big\|\partial_{x}\big(P_{\leq\ell}\phi-P_{\leq m}\psi\big)\big\|_{L^{2}}^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}+2^{-n_{2}}M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
Applying (2.13), we see that
$\big\|\partial_{x}\big(u^{(\ell)}-v^{(m)}\big)\big\|_{L^{\infty}_{T}L^{2}}^{2}$
$\lesssim\Big(d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}\Big(\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}.$
Combining this with (3.18), we come to
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\big[(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})\big]^{\frac{1}{2}}$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}^{\frac{1}{2}}\Big(\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big).$
Now, we choose $n_{2}$ as $2^{n_{2}}=C(1+M_{0}^{2})$ with $C\gg 1$, and impose further smallness $T\ll\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ if necessary, so that the coefficient of $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$ on the right-hand side becomes sufficiently small. Then, this term is absorbed into the left-hand side and we finally obtain
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m})(1+M_{0}^{3}),$
which is the claimed bound. ∎
At this stage, we are now ready to complete the proof of Theorem 1.1.
Proof of Theorem 1.1.
Let $\phi\in\mathcal{Z}^{2}$ satisfy $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$, and let $T_{*}>0$ be the constant given in Proposition 3.3. We divide the proof into several steps.
Step 1. We shall construct a local solution $u\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ to (gINLS) with initial data $\phi$ such that $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$ and
(3.24) $\big\|\mathcal{E}^{2}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim\mathcal{E}^{2}(\phi).$
We also show that if $\phi\in\mathcal{Z}^{k}$ for some integer $k\geq 3$, then $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{k})$ and it satisfies
(3.25) $\big\|\mathcal{E}^{k}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim_{k}\mathcal{E}^{k}(\phi).$
Let $\{u^{(\ell)}\}_{\ell\geq 0}$ be the sequence of solutions to the mollified problem (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ constructed in Proposition 3.1. We recall from Propositions 3.2 and 3.3 that $u^{(\ell)}\in C([0,T_{*}];\mathcal{Z}^{2})$ for all $\ell\geq 0$ and $\{u^{(\ell)}\}_{\ell\geq 0}$ is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1})$. Since $C([0,T_{*}];\mathcal{Z}^{1})$ is a complete metric space, there exists $u\in C([0,T_{*}];\mathcal{Z}^{1})$ such that $u^{(\ell)}\to u$ strongly in $C([0,T_{*}];\mathcal{Z}^{1})$. Moreover, (3.5) shows that the sequence $\{\partial_{x}u^{(\ell)}\}_{\ell\geq 0}$ is bounded in $L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. By the Banach-Alaoglu theorem, there exists a subsequence converging weakly-* to some $\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. Since $\partial_{x}u^{(\ell)}\to\partial_{x}u$ strongly in $C([0,T_{*}];L^{2}(\mathbb{R}))$, we have $\partial_{x}u=\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$, that is, $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{2})$. The bound (3.24) follows from (3.5) and weak-* lower semicontinuity of the $L^{\infty}_{T_{*}}H^{1}$ norm. The claim in the regular case $\phi\in\mathcal{Z}^{k}$ ($k\geq 3$) can be shown in the same way by using (3.6).
It remains to show that the limit $u$ is a solution to (gINLS) with $u(0)=\phi$ and belongs to $C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$. The initial condition is clearly satisfied by the strong convergence in $C([0,T_{*}];\mathcal{Z}^{1})$. To see that $u$ satisfies the integral equation
(3.26) $u(t)=U(t)\phi-i\int_{0}^{t}U(t-s)\mathcal{N}(u(s))\,ds,$
we observe that
$\|P_{\leq\ell}\mathcal{N}(P_{\leq\ell}u^{(\ell)})-\mathcal{N}(u)\|_{L^{\infty}_{T_{*}}L^{2}}$
$\lesssim 2^{-\ell}\|\mathcal{N}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}H^{1}}+\|\mathcal{N}(P_{\leq\ell}u^{(\ell)})-\mathcal{N}(u)\|_{L^{\infty}_{T_{*}}L^{2}}$
$\lesssim 2^{-\ell}(1+\delta^{-1})\|\mathcal{E}^{2}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}^{\frac{3}{2}}+(1+\delta^{-1})\Big(1+\|\mathcal{E}^{1}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}+\|\mathcal{E}^{1}(u)\|_{L^{\infty}_{T_{*}}}\Big)\|d^{1}(P_{\leq\ell}u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{\delta,M_{0}}2^{-\ell}+\|d^{1}(u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
$\to\ 0\qquad(\ell\to\infty),$
where we have used Lemma 2.1, (2.10), (3.5), (3.24) and the strong convergence $u^{(\ell)}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1})$. In particular, we have the convergence of the Duhamel integrals
$i\int_{0}^{t}U(t-s)P_{\leq\ell}\mathcal{N}(P_{\leq\ell}u^{(\ell)}(s))\,ds\ \to\ -i\int_{0}^{t}U(t-s)\mathcal{N}(u(s))\,ds\qquad\text{strongly in $C([0,T_{*}];L^{2}(\mathbb{R}))$}.$
This and the convergence of the linear part imply that $u$ satisfies (3.26). Since $U(t)\phi\in C(\mathbb{R};\mathcal{Z}^{1}_{\rho_{\phi}})$ by (2.17) and $\int_{0}^{t}U(t-s)\mathcal{N}(u(s))\,ds\in C([0,T_{*}];H^{1}(\mathbb{R}))$ for $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ by Lemma 2.1, we see from the integral equation (3.26) that $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$.
We conclude Step 1 by mentioning a bound of distance between two solutions of (gINLS). Let $\phi,\psi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi),\mathcal{E}^{2}(\psi)\leq M_{0}^{2}$, and let $u,v\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ be the corresponding solutions of (gINLS) constructed above. Then, taking $\ell,m\to\infty$ in (3.15), we see that
(3.27) $\big\|d^{1}(u,v)\big\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(\phi,\psi).$
Step 2. We claim the following: Let $u,v\in C([0,T];\mathcal{Z}^{2})$ be two solutions to (gINLS) on an interval $[0,T]$. Assume that $\|\mathcal{E}^{2}(u)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ and $\|\mathcal{E}^{2}(v)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ for some $M_{1}>0$. Then, there exists $T_{1}=T_{1}(\delta,M_{1})>0$ such that for $0\leq t\leq\min\{T,T_{1}\}$ it holds that
$d^{1}(u(t),v(t))\lesssim(1+M_{1})d^{1}(u(0),v(0)).$
In particular, the solution of the Cauchy problem associated with (gINLS) is unique in $C([0,T];\mathcal{Z}^{2})$.
We remark that this claim must be discussed independently of the previous Step 1 (in particular the difference estimate (3.27)), since for uniqueness we have to consider general solutions which may differ from those constructed in Step 1. However, the claim can be proven along the same lines as Proposition 3.3 using the original equation (gINLS), (3.26) instead of the mollified one (3.1), (3.2). We omit the details.
Step 3. We shall prove that the solution $u$ constructed in Step 1 satisfies $u\in C([0,T_{*}];\mathcal{Z}^{2})$ and the continuity of the flow map $u_{0}\mapsto u$ in $C([0,T_{*}];\mathcal{Z}^{2})$.
We can prove $u\in C([0,T_{*}];\mathcal{Z}^{k})$ for $\phi\in\mathcal{Z}^{k}$ if $k\geq 3$ in a similar manner; we omit the proof. In particular, this concludes the proof of Theorem 1.1. To prove the claim, we may adapt an argument due to Bona–Smith [8] or the idea of frequency envelope due to Tao [42, 23, 3] to the non-vanishing solutions in $\mathcal{Z}^{2}$.444 The abstract theorem in [3] does not seem to apply directly to our problem, since $\mathcal{Z}^{2}$ is not a vector space. Here, we outline an adaptation of the latter idea.
Fix $\varepsilon\in(0,1)$, and define the frequency envelope of $f\in\mathcal{Z}^{2}$ by
$c_{j}[f]:=\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\|P_{k}\partial_{x}f\|_{H^{1}}+2^{-\varepsilon j}\big(\|f\|_{L^{\infty}}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}^{\frac{1}{2}}\big),\qquad j\in\mathbb{Z}_{\geq 0}.$
We collect some basic properties of $\mathbf{c}[f]:=\{c_{j}[f]\}_{j\geq 0}$ in the following lemma, which can be shown by a standard argument (see, e.g., [23]); we give a proof in Appendix B for completeness.
Lemma 3.4.
For any $f,g\in\mathcal{Z}^{2}$ and integers $j,k\geq 0$, the following hold.
  1. $\mathrm{(i)}$
    $\|P_{j}\partial_{x}f\|_{H^{1}}\leq c_{j}[f]$, $c_{k}[f]\leq 2^{\varepsilon|k-j|}c_{j}[f]$, $\mathcal{E}^{2}(f)\sim\|\mathbf{c}[f]\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}^{2}$.
  2. $\mathrm{(ii)}$
    $\big\|\mathbf{c}[f]-\mathbf{c}[g]\big\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}\lesssim\|\partial_{x}(f-g)\|_{H^{1}}+\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}}^{1/2}$.
  3. $\mathrm{(iii)}$
    $\mathcal{E}^{3}(P_{\leq j}f)]^{1/2}\lesssim 2^{j}c_{j}[f]$.
For any $\phi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$ and $j\geq 1$, let $u$ and $u^{j}$ be the unique solutions of (gINLS) with initial condition, respectively, $u(0)=\phi$ and $u^{j}(0)=P_{\leq j}\phi$, which we have constructed in the class $C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$.555 From (2.9), we see that $\mathcal{E}^{2}(P_{\leq j}\phi)\leq 3M_{0}^{2}$ for any $j\geq 0$. Therefore, by slightly modifying the choice of $T_{*}$ in Proposition 3.3 we can construct the solutions $u$, $u^{j}$ on a uniform interval $0,T_{*}]$. Note that $P_{\leq j}\phi\in\mathcal{Z}^{3}$ and hence $u^{j}\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{3})\subset C([0,T_{*}];\mathcal{Z}^{2})$. Using the $\mathcal{Z}^{3}$ bound from (3.25), the $\mathcal{Z}^{1}$ difference bound (3.27), and Lemma 3.4, we can derive the following estimate:
(3.28) $\big\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\big\|_{L^{\infty}_{T_{*}}L^{2}}\lesssim_{M_{0}}2^{-|k-j|}c_{j}[\phi],\qquad j\geq 1,\quad k\geq 0.$
This estimate is the key to prove the claim of Step 3 and shows that the initial frequency distribution is not significantly disturbed during the local existence time. For the reader’s convenience, we also give a proof of (3.28) in Appendix B.
To see $u\in C([0,T_{*}];\mathcal{Z}^{2})$, we first notice that the difference bound (3.27) shows the convergence $u^{j}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1})$, and hence $u\in C([0,T_{*}];\mathcal{Z}^{1})$. It then suffices to prove $u\in C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$. We deduce from (3.28) that for any $m_{2}>m_{1}\geq 0$,
$\|\partial_{x}^{2}(u^{m_{2}}-u^{m_{1}})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $=\bigg\|\sum_{j=m_{1}+1}^{m_{2}}\partial_{x}^{2}(u^{j}-u^{j-1})\bigg\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$
$\lesssim\sum_{j_{1},j_{2}=m_{1}+1}^{m_{2}}\sum_{\begin{smallmatrix}k_{1},k_{2}\geq 0\\ |k_{1}-k_{2}|\leq 1\end{smallmatrix}}\|P_{k_{1}}\partial_{x}^{2}(u^{j_{1}}-u^{j_{1}-1})\|_{L^{\infty}_{T_{*}}L^{2}}\|P_{k_{2}}\partial_{x}^{2}(u^{j_{2}}-u^{j_{2}-1})\|_{L^{\infty}_{T_{*}}L^{2}}$
$\lesssim_{M_{0}}\sum_{j_{1},j_{2}=m_{1}+1}^{m_{2}}\sum_{\begin{smallmatrix}k_{1},k_{2}\geq 0\\ |k_{1}-k_{2}|\leq 1\end{smallmatrix}}2^{-|k_{1}-j_{1}|}c_{j_{1}}[\phi]2^{-|k_{2}-j_{2}|}c_{j_{2}}[\phi]$
$\lesssim\sum_{j_{1},j_{2}=m_{1}+1}^{m_{2}}2^{-\frac{1}{2}|j_{1}-j_{2}|}c_{j_{1}}[\phi]c_{j_{2}}[\phi]\lesssim\sum_{j=m_{1}+1}^{m_{2}}c_{j}[\phi]^{2}\leq\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>m_{1}})}^{2}.$
Since $\mathbf{c}[\phi]\in\ell^{2}(\mathbb{Z}_{\geq 0})$, we see that the sequence $\{u^{j}\}_{j\geq 0}$ is Cauchy in $C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$, concluding that the limit $u$ belongs to $C([0,T_{*}];\mathcal{Z}^{2})$ and that
(3.29) $\|\partial_{x}^{2}(u-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}\lesssim_{M_{0}}\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})},\qquad j\geq 0.$
To see the continuity of the flow map, let $\phi,\{\phi_{n}\}_{n\geq 0}$ be such that $\mathcal{E}^{2}(\phi),\mathcal{E}^{2}(\phi_{n})\leq M_{0}^{2}$ and $d^{2}(\phi_{n},\phi)\to 0$ as $n\to\infty$. Furthermore, let $u_{n},u_{n}^{j}\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ denote the solutions to (gINLS) with $u_{n}(0)=\phi_{n}$ and $u_{n}^{j}(0)=P_{\leq j}\phi_{n}$. In view of the $\mathcal{Z}^{1}$ difference estimate (3.27), it suffices to prove
(3.30) $\lim_{n\to\infty}\|\partial_{x}^{2}(u_{n}-u)\|_{L_{T_{*}}^{\infty}L^{2}}=0.$
The triangle inequality yields
$\|\partial_{x}^{2}(u_{n}-u)\|_{L_{T_{*}}^{\infty}L^{2}}\leq\|\partial_{x}^{2}(u_{n}-u_{n}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{2}(u-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{2}(u_{n}^{j}-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}.$
Using (3.29) and Lemma 3.4 (ii), we bound the first two terms on the right-hand side by
$\|\mathbf{c}[\phi_{n}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}\lesssim d^{2}(\phi_{n},\phi)+\big[d^{2}(\phi_{n},\phi)\big]^{\frac{1}{2}}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}.$
For the last term, we can use the $\mathcal{Z}^{3}$ bound from (3.25) and the difference bound (3.27) as well as (2.12):
(3.31) $\|\partial_{x}^{2}(u_{n}^{j}-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\leq\big(\|\partial_{x}^{3}u_{n}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{3}u^{j}\|_{L^{\infty}_{T_{*}}L^{2}}\big)\|d^{1}(u_{n}^{j},u^{j})\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{M_{0}}2^{j}\,d^{1}(P_{\leq j}\phi_{n},P_{\leq j}\phi)\lesssim_{M_{0}}2^{j}\,d^{1}(\phi_{n},\phi).$
Putting these estimates together, we have
$\limsup_{n\to\infty}\|\partial_{x}^{2}(u_{n}-u)\|_{L^{\infty}_{T_{*}}L^{2}}\lesssim_{M_{0}}\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}$
for any $j$, and obtain (3.30) by letting $j\to\infty$.
We have completed the proof of Theorem 1.1. ∎
We conclude this section by showing the convergence result.
Proof of Corollary 1.3.
Let $\{\phi_{\delta}\}_{\delta\in[1,\infty]}$, $\{u_{\delta}\}_{\delta\in[1,\infty]}$ be as in the statement of the corollary, and take $M_{0}>0$ such that $\mathcal{E}^{2}(\phi_{\infty})\leq M_{0}^{2}$. Then, there exists $\delta_{*}\in[1,\infty)$ such that $\mathcal{E}^{2}(\phi_{\delta})\leq 2M_{0}^{2}$ and $\mathcal{E}^{2}(P_{\leq j}\phi_{\delta})\leq 6M_{0}^{2}$ for any $\delta\in[\delta_{*},\infty]$ and $j\in\mathbb{Z}_{\geq 0}$. Let $u_{\delta}^{j}$ denote the (unique) solution to (gINLS) ($\delta\in[1,\infty)$) or (gCM) ($\delta=\infty$) with $u_{\delta}^{j}(0)=P_{\leq j}\phi_{\delta}$. From the proof of Theorem 1.1, there exists $T_{*}\gtrsim(1+M_{0}^{4})^{-1}$ such that for any $\delta\in[\delta_{*},\infty]$ and $j\in\mathbb{Z}_{\geq 0}$, these solutions exist on $0,T_{*}]$ (in particular $T_{\delta,\max}>T_{*}$) and satisfy
(3.32) $\|\mathcal{E}^{2}(u_{\delta})\|_{L^{\infty}_{T_{*}}}+\|\mathcal{E}^{2}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim M_{0}^{2},$
(3.33) $\|\mathcal{E}^{3}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim\mathcal{E}^{3}(P_{\leq j}\phi_{\delta})\lesssim 2^{2j}M_{0}^{2},$
(3.34) $\|d^{1}(u_{\delta}^{j},u_{\delta})\|_{L^{\infty}_{T_{*}}}$ $\lesssim_{M_{0}}d^{1}(P_{\leq j}\phi_{\delta},\phi_{\delta}),$
(3.35) $\|\partial_{x}^{2}(u_{\delta}-u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\lesssim_{M_{0}}\|\mathbf{c}[\phi_{\delta}]\|_{\ell^{2}(\mathbb{Z}_{>j})}.$
Indeed, these estimates have been shown in (3.24), (3.25), (3.27), and (3.29), respectively, with the constants independent of $\delta\in[1,\infty]$.
We next prove the difference estimate
(3.36) $\|d^{1}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}(1+M_{0}),\qquad\delta\in[\delta_{*},\infty).$
To see (3.36), it suffices to prove
(3.37) $\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-\frac{1}{2}}(1+M_{0}),\qquad\delta\in[\delta_{*},\infty),\ j\geq 0$
and take the limit $j\to\infty$ with the estimate (3.34). Then, since $u_{\delta}^{j}$ is smooth, we can follow the calculations in the proof of Proposition 3.3. In fact, the main part of the proof involving the modified energy estimate remains unchanged; we only have to replace the difference of $\mathcal{N}_{\delta}$’s for two solutions by a single $\mathcal{N}_{\delta}(u_{\delta}^{j})$. From Lemma 2.1 and (3.32) we have
$\|\mathcal{N}_{\delta}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}H^{2}}\lesssim\delta^{-1}M_{0}^{3}.$
Then, the previous argument for Proposition 3.3 implies the estimates
$\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim(1+M_{0})d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-1}T_{*}(1+M_{0})M_{0}^{3}+\|\partial_{x}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}},$
$\|\partial_{x}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\lesssim\|\partial_{x}P_{\leq j}(\phi_{\delta}-\phi_{\infty})\|_{L^{2}}^{2}+\delta^{-1}T_{*}(M_{0}^{4}+M_{0}^{6})$
$\quad+\big(T_{*}(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\big)\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}^{2}$
instead of (3.18) and (3.23), respectively. Choosing $n_{2}$ appropriately and modifying $T_{*}\sim(1+M_{0}^{4})^{-1}$ if necessary, we obtain (3.37).
In view of (3.36), the convergence $\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}\to 0$ follows once $\|\partial_{x}^{2}(u_{\delta}-u_{\infty})\|_{L^{\infty}_{T_{*}}L^{2}}\to 0$ is verified. To prove the latter convergence, we follow the argument showing continuous dependence in Step 3 of the proof of Theorem 1.1. Using interpolation, (3.33), and (3.37) in place of (3.27), we have the fixed-$j$ difference estimate corresponding to (3.31):
$\|\partial_{x}^{2}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\leq\big(\|\partial_{x}^{3}u_{\delta}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{3}u_{\infty}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}\big)\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{M_{0}}2^{j}\big(d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-\frac{1}{2}}\big)$
$\lesssim_{M_{0}}2^{j}\big(d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}\big).$
Combining it with (3.35), we obtain
$\limsup_{\delta\to\infty}\|\partial_{x}^{2}(u_{\delta}-u_{\infty})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\leq\limsup_{\delta\to\infty}\Big(\|\partial_{x}^{2}(u_{\delta}-u_{\delta}^{j})\|_{L^{\infty}_{T*}L^{2}}+\|\partial_{x}^{2}(u_{\infty}-u_{\infty}^{j})\|_{L^{\infty}_{T*}L^{2}}+\|\partial_{x}^{2}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}\Big)$
$\lesssim_{M_{0}}\limsup_{\delta\to\infty}\Big(\|\mathbf{c}[\phi_{\delta}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\|\mathbf{c}[\phi_{\infty}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\big[2^{j}\big(d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}\big)\big]^{\frac{1}{2}}\Big)$
$\lesssim\|\mathbf{c}[\phi_{\infty}]\|_{\ell^{2}(\mathbb{Z}_{>j})}\ \to 0\qquad(j\to\infty).$
We have thus proven that $\lim\limits_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}=0$.
Now, we take $T\in(0,T_{\infty,\max})$ arbitrarily. Since $\mathcal{E}^{2}(u_{\infty}(t))$ is bounded on $[0,T]$, repeating the above procedure finitely many times we see that there exists $\delta_{**}\in[1,\infty)$ such that $T_{\delta,\max}>T$ for any $\delta\in[\delta_{**},\infty]$ and $\lim\limits_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T}}=0$. This completes the proof of Corollary 1.3. ∎
4. Global well-posedness
Our goal in this section is to establish Theorem 1.4. To do so, we derive the conservation laws for (gINLS) in the $\mathcal{Z}^{2}$ framework; in particular, they are meaningful for solutions with nonvanishing boundary conditions at infinity. We then show that, for any real parameters $\alpha\neq 0$, $\beta\geq 0$, the Cauchy problem for (gINLS) is globally well-posed in $\mathcal{Z}^{2}_{*}$. In addition, when $\alpha=\beta=1$ we obtain a uniform bound for global solutions to (INLS). Our argument also yields the corresponding results for (gCM) and (CM-defoc).
The main result in this section is the following:
Theorem 4.1.
Let $\delta>0$, $\alpha,\beta\in\mathbb{R}$. Let $\phi\in\mathcal{Z}^{2}_{\rho}$ for some $\rho\geq 0$, and let $u\in C([0,T_{\max});\mathcal{Z}^{2}_{\rho})$ be the unique maximal solution to the Cauchy problem for (gINLS) with the initial condition $u(0)=\phi$ constructed in Theorem 1.1.
  1. $\mathrm{(i)}$
    The quantity
    $\mathcal{H}_{1}(u)=\int_{\mathbb{R}}\Big\{|\partial_{x}u|^{2}$ $+\alpha(|u|^{2}-\rho^{2})\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{\beta}{2}(|u|^{2}-\rho^{2})T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})$
    $+\alpha^{2}\Big(\frac{1}{3}|u|^{2}+\frac{1}{6}\rho^{2}\Big)(|u|^{2}-\rho^{2})^{2}\Big\}\,dx$
    is well-defined in $\mathcal{Z}^{2}_{\rho}$, and $\mathcal{H}_{1}(u(t))$ is conserved in time.
  2. $\mathrm{(ii)}$
    The quantity
    $\mathcal{H}_{2}(u)=\int_{\mathbb{R}}\Big\{|\partial_{x}^{2}u|^{2}$ $+\alpha\Big(2|\partial_{x}u|^{2}-3\partial_{x}^{2}(|u|^{2})\Big)\operatorname{Im}[\overline{u}\partial_{x}u]$
    $+\beta\Big(2|\partial_{x}u|^{2}-\frac{3}{4}\partial_{x}^{2}(|u|^{2})\Big)T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})$
    $+\beta\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\Big\}\,dx$
    is well-defined in $\mathcal{Z}^{2}_{\rho}$, and $\mathcal{H}_{2}(u(t))$ satisfies, for $t\in[0,T_{\max})$,
    (4.1) $\mathcal{H}_{2}(u(t))\lesssim_{\alpha,\beta,\delta}\mathcal{H}_{2}(u_{0})+\int_{0}^{t}\Big(\mathcal{E}^{1}(u(s))+\big[\mathcal{E}^{1}(u(s))\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u(s)\|_{L^{2}}^{2}\Big)\,ds.$
  3. $\mathrm{(iii)}$
    Assume $\alpha\neq 0$, $\beta\geq 0$ and $\rho>0$. Then, the solution $u$ obeys the following a priori bounds:
    (4.2) $\mathcal{E}^{1}(u(t))$ $\leq C_{1},$
    (4.3) $\|\partial_{x}^{2}u(t)\|_{L^{2}}^{2}$ $\leq C_{2}e^{C_{3}t}(1+\|\partial_{x}^{2}\phi\|_{L^{2}}^{2})$
    for $t\in[0,T_{\max})$, where the constants $C_{1},C_{2},C_{3}>0$ depend only on $\alpha,\beta,\delta,\rho$ and $\mathcal{E}^{1}(\phi)$. In particular, $T_{\max}=\infty$.
Furthermore, for (gCM) we have the same results as above, with $T_{\delta}$ replaced by $H$ and with the constants being independent of $\delta$.
To derive $\mathcal{H}_{1}(u)$ and $\mathcal{H}_{2}(u),$ we proceed by constructing modified energies through a cancellation method [40, 13]. We begin with the time derivative of the leading term (either $\int|\partial_{x}u|^{2}dx$ or $\int|\partial_{x}^{2}u|^{2}dx$) and systematically introduce correctors to cancel out the problematic nonlinear terms. For $\mathcal{H}_{1}(u)$, this process is continued to provide a conserved quantity for general $\alpha,\beta$. Although this is not the case for $\mathcal{H}_{2}(u)$ in general, in the integrable case we can also derive an exact conserved quantity as in the following corollary.
Corollary 4.2.
Let $\delta>0$ and $\phi\in\mathcal{Z}^{2}_{\rho}$ for some $\rho\geq 0$. If $\beta=\pm|\alpha|$, the quantity
$\mathcal{H}_{2}^{\mathrm{INLS}}(u):=\int_{\mathbb{R}}\Big\{\begin{aligned} &|\partial_{x}^{2}u|^{2}+\alpha\Big(2|\partial_{x}u|^{2}-3\partial_{x}^{2}(|u|^{2})\Big)\operatorname{Im}[\overline{u}\partial_{x}u]\\ &+\beta\Big(2|\partial_{x}u|^{2}-\frac{3}{4}\partial_{x}^{2}(|u|^{2})\Big)T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})+\beta\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\\ &+2\alpha^{2}|u|^{4}|\partial_{x}u|^{2}+\alpha^{2}|u|^{2}\Big(\frac{3}{2}[\partial_{x}(|u|^{2})]^{2}+\frac{1}{2}[T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})]^{2}\Big)\\ &+\alpha\beta\operatorname{Im}[\overline{u}\partial_{x}u]\Big(2|u|^{2}T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})+T_{\delta}\partial_{x}(|u|^{4}-\rho^{4})\Big)\\ &+\alpha^{2}(|u|^{6}-\rho^{6})\Big(\alpha\operatorname{Im}[\overline{u}\partial_{x}u]+\frac{2}{3}\beta T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})\Big)+\frac{1}{4}\alpha^{2}\beta(|u|^{4}-\rho^{4})T_{\delta}\partial_{x}(|u|^{4}-\rho^{4})\\ &+\alpha^{4}\Big(\frac{1}{5}|u|^{6}+\frac{2}{5}\rho^{2}|u|^{4}+\frac{3}{5}\rho^{4}|u|^{2}+\frac{3}{10}\rho^{6}\Big)(|u|^{2}-\rho^{2})^{2}\Big\}\,dx\end{aligned}$
is well-defined in $\mathcal{Z}^{2}_{\rho}$ and is a conserved quantity for (gINLS). If $\beta=|\alpha|>0$ and $\rho>0$, the global solution $u\in C(\mathbb{R};\mathcal{Z}^{2}_{\rho})$ to (gINLS) satisfies
(4.4) $\sup_{t\in\mathbb{R}}\mathcal{E}^{2}(u(t))<\infty.$
Furthermore, for (gCM) we have the same results as above, with $T_{\delta}$ replaced by $H$.
Remark 4.1.
For vanishing solutions (in $H^{s}$), infinitely many conservation laws for (CM-defoc), (CM-foc) and for (INLS) have been obtained in [25, 19] and in [9], respectively. In particular, a conserved quantity at the $H^{k}(\mathbb{R})$ level is given by $\|\mathcal{L}_{u(t)}^{k}u(t)\|_{L^{2}}^{2}$, where $\mathcal{L}_{u}$ is a suitable Lax operator (see (1.5) for the case of (INLS)). In [11, Lemma 3.5], Chen found conserved quantities at the $\mathcal{Z}^{1}$ and $\mathcal{Z}^{2}$ levels for (CM-defoc), which were obtained by expanding $\|\mathcal{L}_{u}u\|_{L^{2}}$, $\|\mathcal{L}_{u}^{2}u\|_{L^{2}}$ and “renormalizing” them for nonvanishing solutions, i.e., replacing $\Pi(|u|^{2})$ with $\Pi(|u|^{2}-1)+1$, where $\Pi=\frac{1}{2}(1+iH)$ is the Cauchy-Szegő projection and is a bounded operator on $L^{2}(\mathbb{R})$. It is then natural to ask whether conservation laws for nonvanishing solutions to (INLS) can be obtained by similar renormalization procedure. However, for (INLS) the operator $\frac{1}{2}(1+iT_{\delta})$ corresponding to $\Pi$ becomes unbounded on $L^{2}(\mathbb{R})$, and there is no apparent way of renormalizing $\|\mathcal{L}_{u}^{k}u\|_{L^{2}}$ for $u\in\mathcal{Z}^{k}_{\rho}$. Instead, in Theorem 4.1 and Corollary 4.2, we construct conserved quantities directly by using the idea of modified energy method.
From now on, we only consider (gINLS); however, we can apply exactly the same argument to (gCM) by replacing $T_{\delta}$ with $H$. Before proving Theorem 4.1, we collect some useful lemmas.
Lemma 4.1.
Let $u\in C([0,T];\mathcal{Z}^{\infty})$ be a smooth solution of (gINLS). Then,
$\partial_{t}(|u|^{2})$ $=\partial_{x}\Big(2\operatorname{Im}[\overline{u}\partial_{x}u]+\alpha|u|^{4}\Big),$
$\partial_{t}(|\partial_{x}u|^{2})$ $=\partial_{x}\Big(2\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]+\frac{\alpha}{2}[\partial_{x}(|u|^{2})]^{2}\Big)+2\alpha|\partial_{x}u|^{2}\partial_{x}(|u|^{2})+2\beta\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}^{2}(|u|^{2}),$
$\partial_{t}(|\partial_{x}^{2}u|^{2})$ $=\partial_{x}\Big(2\operatorname{Im}[(\partial_{x}^{2}\overline{u})\partial_{x}^{3}u]+\frac{\alpha}{2}[\partial_{x}^{2}(|u|^{2})]^{2}\Big)$
$\quad+2\alpha|\partial_{x}^{2}u|^{2}\partial_{x}(|u|^{2})+2\alpha\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2})-2\alpha|\partial_{x}u|^{2}\partial_{x}^{3}(|u|^{2})$
$\quad+4\beta\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]T_{\delta}\partial_{x}^{2}(|u|^{2})+2\beta\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\cdot T_{\delta}\partial_{x}^{3}(|u|^{2}),$
$\partial_{t}(\operatorname{Im}[\overline{u}\partial_{x}u])$ $=2\partial_{x}(|\partial_{x}u|^{2})-\frac{1}{2}\partial_{x}^{3}(|u|^{2})+2\alpha\operatorname{Im}[\overline{u}\partial_{x}u]\partial_{x}(|u|^{2})+\beta|u|^{2}T_{\delta}\partial_{x}^{2}(|u|^{2}).$
Proof.
The proof follows from a direct calculation. ∎
Lemma 4.2.
Let $u\in L^{\infty}(\mathbb{R})$ satisfy $\partial_{x}u\in H^{1}(\mathbb{R})$. Then, we have
$\int\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])=\frac{1}{2}\int|\partial_{x}^{2}u|^{2}\partial_{x}(|u|^{2})-\frac{1}{4}\int\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2}).$
Proof.
Using the identity $\operatorname{Im}(z_{1})\operatorname{Im}(z_{2})=\operatorname{Re}(z_{1}\overline{z_{2}})-\operatorname{Re}(z_{1})\operatorname{Re}(z_{2})$ ($z_{1},z_{2}\in\mathbb{C}$), we have
$\int\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])=\int|\partial_{x}^{2}u|^{2}\operatorname{Re}[u\partial_{x}\overline{u}]-\int\operatorname{Re}[(\partial_{x}\overline{u})\partial_{x}^{2}u]\operatorname{Re}[\overline{u}\partial_{x}^{2}u]$
$=\int|\partial_{x}^{2}u|^{2}\cdot\frac{1}{2}\partial_{x}(|u|^{2})-\int\frac{1}{2}\partial_{x}(|\partial_{x}u|^{2})\cdot\Big(\frac{1}{2}\partial_{x}^{2}(|u|^{2})-|\partial_{x}u|^{2}\Big)$
$=\frac{1}{2}\int|\partial_{x}^{2}u|^{2}\partial_{x}(|u|^{2})-\frac{1}{4}\int\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2}),$
as desired. ∎
We are now in a position to prove Theorem 4.1.
Proof of Theorem 4.1.
First, it is easy to see that $\mathcal{H}_{1}(u)$ and $\mathcal{H}_{2}(u)$ are well-defined for $u\in\mathcal{Z}^{2}_{\rho}$, noticing $u\in L^{\infty}(\mathbb{R})$, $\partial_{x}u\in H^{1}(\mathbb{R})$, $|u|^{2}-\rho^{2}\in H^{1}(\mathbb{R})$, and that $T_{\delta}\partial_{x}$ maps $H^{1}(\mathbb{R})$ into $L^{2}(\mathbb{R})$ by (2.3). Moreover, we may assume that $u\in\mathcal{Z}^{\infty}_{\rho}$ by a standard approximation argument, which relies on continuity of the flow map and persistence of regularity in the local well-posedness Theorem 1.1.
For (i), we begin by observing, from Lemma 4.1, that
$\partial_{t}\int|\partial_{x}u|^{2}=2\alpha\int|\partial_{x}u|^{2}\partial_{x}(|u|^{2})+2\beta\int\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}^{2}(|u|^{2}).$
To cancel out two terms on the right-hand side, we introduce two correctors
$I_{1}^{(1)}(u):=\int(|u|^{2}-\rho^{2})\operatorname{Im}[\overline{u}\partial_{x}u],\qquad I_{1}^{(2)}(u):=\int(|u|^{2}-\rho^{2})T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}).$
Using Lemma 4.1 and integration by parts, we see that
$\partial_{t}I_{1}^{(1)}(u)$ $=-2\int|\partial_{x}u|^{2}\partial_{x}(|u|^{2})-2\alpha\int|u|^{2}(|u|^{2}-\rho^{2})\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])$
$\quad-\beta\int\partial_{x}(|u|^{4})\cdot T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}),$
$\partial_{t}I_{1}^{(2)}(u)$ $=-4\int\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}^{2}(|u|^{2})+2\alpha\int\partial_{x}(|u|^{4})\cdot T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}).$
Here, we have used the symmetry of $T_{\delta}\partial_{x}$ (2.5) to verify
$\int\partial_{x}(|u|^{2})\cdot T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})=\int\partial_{x}(|u|^{2}-\rho^{2})\cdot T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})=0.$
From these calculations, we see that
$\partial_{t}\Big(\int|\partial_{x}u|^{2}+\alpha I_{1}^{(1)}(u)+\frac{\beta}{2}I_{1}^{(2)}(u)\Big)=-2\alpha^{2}\int|u|^{2}(|u|^{2}-\rho^{2})\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]).$
To cancel out the right-hand side, we further add two correctors
$I_{1}^{(3)}(u):=\int(|u|^{2}-\rho^{2})^{3},\qquad I_{1}^{(4)}(u):=\int(|u|^{2}-\rho^{2})^{2}.$
Using Lemma 4.1 and integration by parts, we see that
$\partial_{t}I_{1}^{(3)}(u)$ $=6\int(|u|^{2}-\rho^{2})^{2}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]),$
$\partial_{t}I_{1}^{(4)}(u)$ $=4\int(|u|^{2}-\rho^{2})\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]).$
Here, we have calculated as follows: for $k\geq 1$,
$\int(|u|^{2}-\rho^{2})^{k}\partial_{x}(|u|^{4})$ $=\int(|u|^{2}-\rho^{2})^{k}\partial_{x}\Big((|u|^{2}-\rho^{2})^{2}+2\rho^{2}(|u|^{2}-\rho^{2})+\rho^{4}\Big)$
$=\int\partial_{x}\Big(\frac{2}{k+2}(|u|^{2}-\rho^{2})^{k+2}+\frac{2\rho^{2}}{k+1}(|u|^{2}-\rho^{2})^{k+1}\Big)=0.$
From the above calculations, we have
$2\alpha^{2}\partial_{t}\Big(\frac{1}{6}I_{1}^{(3)}(u)+\frac{\rho^{2}}{4}I_{1}^{(4)}(u)\Big)=2\alpha^{2}\int|u|^{2}(|u|^{2}-\rho^{2})\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]).$
Consequently, the quantity
$\mathcal{H}_{1}(u):=\int|\partial_{x}u|^{2}+\alpha I_{1}^{(1)}(u)+\frac{\beta}{2}I_{1}^{(2)}(u)+\frac{\alpha^{2}}{3}I_{1}^{(3)}(u)+\frac{\alpha^{2}\rho^{2}}{2}I_{1}^{(4)}(u)$
is conserved in time. This finishes the proof of (i).
In a similar manner, to compute $\mathcal{H}_{2}$, we begin with
$\partial_{t}\int|\partial_{x}^{2}u|^{2}$ $=2\alpha\int|\partial_{x}^{2}u|^{2}\partial_{x}(|u|^{2})+4\alpha\int\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2})$
$\quad+4\beta\int\operatorname{Im}[(\partial_{x}\overline{u})\partial_{x}^{2}u]T_{\delta}\partial_{x}^{2}(|u|^{2})+2\beta\int\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\cdot T_{\delta}\partial_{x}^{3}(|u|^{2}),$
and to cancel out the terms on the right-hand side we introduce five correctors666 We can predict these terms from the terms of degree four in $E_{2}(u)$ given in Section 1.
$I_{2}^{(1)}(u):=\int|\partial_{x}u|^{2}\operatorname{Im}[\overline{u}\partial_{x}u],\qquad I_{2}^{(2)}(u):=\int\partial_{x}^{2}(|u|^{2})\cdot\operatorname{Im}[\overline{u}\partial_{x}u],$
$I_{2}^{(3)}(u):=\int|\partial_{x}u|^{2}T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}),\qquad I_{2}^{(4)}(u):=\int\partial_{x}^{2}(|u|^{2})\cdot T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}),$
$I_{2}^{(5)}(u):=\int\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]).$
By a calculation involving Lemmas 4.14.2, we see that
$\partial_{t}I_{2}^{(1)}(u)$ $=-\int|\partial_{x}^{2}u|^{2}\partial_{x}(|u|^{2})+\int\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2})\ +R_{1}(u),$
$\partial_{t}I_{2}^{(2)}(u)$ $=2\int\partial_{x}(|\partial_{x}u|^{2})\cdot\partial_{x}^{2}(|u|^{2})\ +R_{2}(u),$
$\partial_{t}I_{2}^{(3)}(u)$ $=-2\int\operatorname{Im}[\partial_{x}\overline{u}\cdot\partial_{x}^{2}u]T_{\delta}\partial_{x}^{2}(|u|^{2})-2\int\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}^{2}(|\partial_{x}u|^{2})\ +R_{3}(u),$
$\partial_{t}I_{2}^{(4)}(u)$ $=4\int\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\cdot T_{\delta}\partial_{x}^{3}(|u|^{2})\ +R_{4}(u),$
$\partial_{t}I_{2}^{(5)}(u)$ $=\int\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\cdot T_{\delta}\partial_{x}^{3}(|u|^{2})+4\int\operatorname{Im}[\overline{u}\partial_{x}u]T_{\delta}\partial_{x}^{2}(|\partial_{x}u|^{2})\ +R_{5}(u).$
Here, each of the remainder terms $R_{j}(u)$ ($j=1,\dots,5$) is a linear combination of integrals with six $u$’s and four spatial differentials, and also of the form
$\text{either}\qquad\int fh_{1}h_{2}\qquad\text{or}\qquad\int g_{1}g_{2}h$
with
$f\in\big\{1,|u|^{2}\big\},\qquad g_{1},g_{2}\in\big\{\partial_{x}(|u|^{2}),\operatorname{Im}[\overline{u}\partial_{x}u],T_{\delta}\partial_{x}(|u|^{2})\big\},$
$h,h_{1},h_{2}\in\big\{|\partial_{x}u|^{2},\partial_{x}^{2}(|u|^{2}),\partial_{x}^{2}(|u|^{4}),\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u]),T_{\delta}\partial_{x}^{2}(|u|^{2}),T_{\delta}\partial_{x}^{2}(|u|^{4}),T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\big\}.$
By using Hölder, Gagliardo-Nirenberg $\|\partial_{x}u\|_{L^{\infty}}\leq\|\partial_{x}u\|_{L^{2}}^{1/2}\|\partial_{x}^{2}u\|_{L^{2}}^{1/2}$, and the estimate (2.3), we estimate these remainder terms as
$|R_{j}(u)|$ $\lesssim_{\alpha,\beta,\delta}\Big(\|u\|_{L^{\infty}}+\|\partial_{x}u\|_{L^{2}}\Big)^{4}\Big(\|u\|_{L^{\infty}}^{2}+\|\partial_{x}u\|_{L^{2}}^{2}+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big)$
$\lesssim\Big(\mathcal{E}^{1}(u)+\big[\mathcal{E}^{1}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
We also see that
$\mathcal{H}_{2}(u):=\int|\partial_{x}^{2}u|^{2}+2\alpha I_{2}^{(1)}(u)-3\alpha I_{2}^{(2)}(u)+2\beta I_{2}^{(3)}(u)-\frac{3}{4}\beta I_{2}^{(4)}(u)+\beta I_{2}^{(5)}(u)$
is the only choice of the coefficients to cancel out the right-hand side of $\partial_{t}\int|\partial_{x}^{2}u|^{2}$ and yield
$\partial_{t}|\mathcal{H}_{2}(u)|\lesssim_{\alpha,\beta}\sum_{j=1}^{5}|R_{j}(u)|\lesssim_{\alpha,\beta,\delta}\Big(\mathcal{E}^{1}(u)+\big[\mathcal{E}^{1}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
Integrating in $t$, we obtain (4.1) and thus (ii).
It remains to prove (iii). By Young’s inequality, we see that
$\Big|\alpha\int(|u|^{2}-\rho^{2})\operatorname{Im}[\overline{u}\partial_{x}u]\Big|\leq\frac{5}{6}\int|\partial_{x}u|^{2}+\frac{3}{10}\alpha^{2}\int|u|^{2}(|u|^{2}-\rho^{2})^{2}.$
Thus, noticing the positivity of $T_{\delta}\partial_{x}$ in (2.5) and the assumption $\beta\geq 0$, we see that
$\mathcal{H}_{1}(u)\geq\frac{1}{6}\|\partial_{x}u\|_{L^{2}}^{2}+\frac{\alpha^{2}\rho^{2}}{6}\||u|^{2}-\rho^{2}\|_{L^{2}}^{2}.$
Using the inequality (2.6) and that $\alpha^{2}\rho^{2}>0$, we conclude
(4.5) $\|\partial_{x}u\|_{L^{2}}^{2}\lesssim\mathcal{H}_{1}(u),\quad\||u|^{2}-\rho^{2}\|_{L^{2}}\lesssim_{\alpha,\rho}\mathcal{H}_{1}(u)^{\frac{1}{2}},\quad\|u\|_{L^{\infty}}^{2}\lesssim_{\alpha,\rho}\mathcal{H}_{1}(u)^{\frac{2}{3}}+1.$
On the other hand, using (2.3)–(2.4) we easily estimate $\mathcal{H}_{1}(\phi)$ as
(4.6) $\mathcal{H}_{1}(\phi)$ $\lesssim_{\alpha,\beta,\rho,\delta}\|\partial_{x}\phi\|_{L^{2}}^{2}+\||\phi|^{2}-\rho^{2}\|_{L^{2}}\|\phi\|_{L^{\infty}}\|\partial_{x}\phi\|_{L^{2}}+\Big(\|\phi\|_{L^{\infty}}^{2}+1\Big)\||\phi|^{2}-\rho^{2}\|_{L^{2}}^{2}$
$\lesssim\big[\mathcal{E}^{1}(\phi)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}(\phi)\big]^{\frac{3}{2}}.$
From (4.5), (4.6) and the conservation of $\mathcal{H}_{1}(u(t))$ in (i), we obtain (4.2).
Finally, we prove (4.3). Using (2.3)–(2.4), Hölder’s inequality and the a priori bound (4.2), we see that
$\|\partial_{x}^{2}(|u|^{2})\|_{L^{2}}+\||\partial_{x}u|^{2}\|_{L^{2}}+\|T_{\delta}\partial_{x}(\operatorname{Im}[\overline{u}\partial_{x}u])\|_{L^{2}}\lesssim 1+\|\partial_{x}^{2}u\|_{L^{2}},$
$\|\operatorname{Im}[\overline{u}\partial_{x}u]\|_{L^{2}}+\|T_{\delta}\partial_{x}(|u|^{2})\|_{L^{2}}\lesssim 1,$
and then, by Young’s inequality, that
$\Big|\mathcal{H}_{2}(u)-\int|\partial_{x}^{2}u|^{2}\Big|\leq\frac{1}{2}\int|\partial_{x}^{2}u|^{2}+C_{0},$
where $C_{0}>0$ and other implicit constants may depend on $\alpha,\beta,\delta,\rho,\mathcal{E}^{1}(\phi)$. In particular, we have
(4.7) $\int|\partial_{x}^{2}u|^{2}\leq 2\mathcal{H}_{2}(u)+2C_{0},\qquad\mathcal{H}_{2}(u)\leq\frac{3}{2}\int|\partial_{x}^{2}u|^{2}+C_{0}.$
The desired a priori bound (4.3) follows from (4.1), (4.2), (4.7) and Gronwall’s inequality.
This is the end of the proof of Theorem 4.1. ∎
At the end of this section, we give an outline of the proof of Corollary 4.2.
Outline of proof of Corollary 4.2.
In the integrable case of (gINLS) where $\beta=\pm|\alpha|$, we can further repeat the energy correction process (as in the proof of Theorem 4.1 (ii)) to obtain an exactly conserved quantity $\mathcal{H}^{\mathrm{INLS}}_{2}(u)$, by introducing the following new correctors of order six with two spatial differentials
$I_{2}^{(6)}(u):=\int|u|^{4}|u_{x}|^{2},\qquad I_{2}^{(7)}(u):=\int|u|^{2}\big[(|u|^{2})_{x}\big]^{2},$
$I_{2}^{(8)}(u):=\int|u|^{2}\operatorname{Im}[\overline{u}u_{x}]T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}),\qquad I_{2}^{(9)}(u):=\int\operatorname{Im}[\overline{u}u_{x}]T_{\delta}\partial_{x}(|u|^{4}-\rho^{4}),$
$I_{2}^{(10)}(u):=\int|u|^{2}\big[T_{\delta}\partial_{x}(|u|^{2}-\rho^{2})\big]^{2},$
the correctors of order eight with one differential
$I_{2}^{(11)}(u):=\int(|u|^{6}-\rho^{6})\operatorname{Im}[\overline{u}u_{x}],$
$I_{2}^{(12)}(u):=\int(|u|^{6}-\rho^{6})T_{\delta}\partial_{x}(|u|^{2}-\rho^{2}),\qquad I_{2}^{(13)}(u):=\int(|u|^{4}-\rho^{4})T_{\delta}\partial_{x}(|u|^{4}-\rho^{4}),$
and finally, the corrector of order ten with no differential
$I_{2}^{(14)}(u):=D_{1}\int(|u|^{2}-\rho^{2})^{5}+D_{2}\rho^{2}\int(|u|^{2}-\rho^{2})^{4}+D_{3}\rho^{4}\int(|u|^{2}-\rho^{2})^{3}+D_{4}\rho^{6}\int(|u|^{2}-\rho^{2})^{2}$
$(D_{1},\dots,D_{4}\in\mathbb{R}),$
and successively determining the coefficients to cancel the remainder in (4).
Second, to prove (4.4), we recall that in the proof of (4.3) in Theorem 4.1, we have already seen that
$\sum_{j=1}^{5}|I_{2}^{(j)}(u)|\lesssim 1+\|\partial_{x}^{2}u\|_{L^{2}}.$
Similarly, using Hölder’s inequality and (2.3), we see that
$\sum_{j=6}^{14}|I_{2}^{(j)}(u)|\lesssim 1.$
Here, the implicit constants depend on $\alpha,\delta,\rho$ and $\mathcal{E}^{1}(u)$. As we did for (4.7), by Young’s inequality we deduce
$\Big|\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))-\int|\partial_{x}^{2}u(t)|^{2}\Big|\leq\frac{1}{2}\int|\partial_{x}^{2}u(t)|^{2}+C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(u(t))\big)\qquad(t\in[0,T])$
for some constant $C_{4}>0$. Then, from the a priori bound (4.2) for $\mathcal{E}^{1}(u(t))$ proven in Theorem 4.1 (here we use the defocusing assumption $\beta=|\alpha|>0$) and conservation of $\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))$, we infer that
$\|\partial_{x}^{2}u(t)\|_{L^{2}}^{2}$ $\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(u(t))\big)$
$\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(\phi)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi)\big)$
$\leq 3\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi)\big)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi))\big),$
where $C_{1}>0$ is the constant in (4.2). This completes the proof. ∎
Appendix A Fourier transform of the hyperbolic cotangent function
We give a proof of (1.1). To this end, we will use the following lemma.
Lemma A.1.
We have
(A.1) $\mathcal{F}\bigg[\frac{1}{e^{x}+1}-\frac{1}{e^{-x}+1}\bigg](\xi)=2i\pi\,\textrm{p.v.}\operatorname{cosech}{(\pi\xi)},$
where p.v. denotes the principal value.
Proof of Lemma A.1.
Let $0<\varepsilon<1/2$ be a constant to be taken $\varepsilon\to 0$. We consider the perimeter of the rectangle $\{z\in\mathbb{C}\colon-R\leq\operatorname{Re}[z]\leq R,~0\leq\operatorname{Im}[z]\leq 2\pi\}$. Then, an elementary complex analysis involving the residue theorem shows that
$\mathcal{F}\left[\frac{e^{\varepsilon x}}{e^{x}+1}\right](\xi)$ $=\int_{-\infty}^{\infty}\frac{e^{(\varepsilon-i\xi)x}}{e^{x}+1}\,dx$
$=2i\pi\,\mathrm{Res}\left(\frac{e^{(\varepsilon-i\xi)z}}{e^{z}+1};i\pi\right)+\lim_{R\to\infty}\int_{-R}^{R}\frac{e^{(\varepsilon-i\xi)(x+2i\pi)}}{e^{x+2i\pi}+1}\,dx$
$\quad-\lim_{R\to\infty}\int_{0}^{2\pi}\frac{e^{(\varepsilon-i\xi)(R+iy)}}{e^{R+iy}+1}i\,dy+\lim_{R\to\infty}\int_{0}^{2\pi}\frac{e^{(\varepsilon-i\xi)(-R+iy)}}{e^{-R+iy}+1}i\,dy$
$=-2i\pi e^{\pi(\xi+i\varepsilon)}+e^{2\pi(\xi+i\varepsilon)}\int_{-\infty}^{\infty}\frac{e^{(\varepsilon-i\xi)x}}{e^{x}+1}\,dx$
$=-2i\pi e^{\pi(\xi+i\varepsilon)}+e^{2\pi(\xi+i\varepsilon)}\mathcal{F}\left[\frac{e^{\varepsilon x}}{e^{x}+1}\right](\xi),$
which implies that
(A.2) $\mathcal{F}\left[\frac{e^{\varepsilon x}}{e^{x}+1}\right](\xi)=\frac{-2i\pi e^{\pi(\xi+i\varepsilon)}}{1-e^{2\pi(\xi+i\varepsilon)}}=\frac{2i\pi}{e^{\pi(\xi+i\varepsilon)}-e^{-\pi(\xi+i\varepsilon)}}.$
Moreover, by $\mathcal{F}[f(-x)](\xi)=\mathcal{F}[f](-\xi)$, we see that
(A.3) $\mathcal{F}\left[\frac{e^{-\varepsilon x}}{e^{-x}+1}\right](\xi)=\mathcal{F}\left[\frac{e^{\varepsilon x}}{e^{x}+1}\right](-\xi)=\frac{2i\pi}{e^{-\pi(\xi-i\varepsilon)}-e^{\pi(\xi-i\varepsilon)}}.$
By (A.2), (A.3), and
$\lim_{\varepsilon\downarrow 0}\left(\frac{e^{\varepsilon x}}{e^{x}+1}-\frac{e^{-\varepsilon x}}{e^{-x}+1}\right)=\frac{1}{e^{x}+1}-\frac{1}{e^{-x}+1}\qquad\mbox{for all $x\in\mathbb{R}$},$
it suffices for (A.1) to show that
(A.4) $F_{\varepsilon}(\xi)$ $=\frac{1}{2i\pi}\mathcal{F}\left[\frac{e^{\varepsilon x}}{e^{x}+1}-\frac{e^{-\varepsilon x}}{e^{-x}+1}\right](\xi)$
$\;=\frac{1}{e^{i\pi\varepsilon}e^{\pi\xi}-e^{-i\pi\varepsilon}e^{-\pi\xi}}-\frac{1}{e^{i\pi\varepsilon}e^{-\pi\xi}-e^{-i\pi\varepsilon}e^{\pi\xi}}$
$\;\to\textrm{p.v.}\operatorname{cosech}{(\pi\xi)}\quad\mbox{in the tempered distribution sense, as $\varepsilon\downarrow 0$}.$
Let $\phi$ be a Schwartz function on $\mathbb{R}$, and $0<r\leq 1$ a constant to be taken $r\to 0$. Since $F_{\varepsilon}$ is an odd function, we see that
(A.5) $\int_{\mathbb{R}}F_{\varepsilon}(\xi)\phi(\xi)\,d\xi=\int_{|\xi|>r}F_{\varepsilon}(\xi)\phi(\xi)\,d\xi+\int_{-r}^{r}F_{\varepsilon}(\xi)\left\{\phi(\xi)-\phi(0)\right\}\,d\xi.$
Note that $F_{\varepsilon}(\xi)$ converges to $\operatorname{cosech}{(\pi\xi)}$ for all $x\in\mathbb{R}\setminus\{0\}$ as $\varepsilon\downarrow 0$. Moreover, we see that
(A.6) $\left|F_{\varepsilon}(\xi)\right|$ $=\left|\frac{1}{e^{i\pi\varepsilon}e^{\pi\xi}-e^{-i\pi\varepsilon}e^{-\pi\xi}}-\frac{1}{e^{i\pi\varepsilon}e^{-\pi\xi}-e^{-i\pi\varepsilon}e^{\pi\xi}}\right|$
$=\left|\frac{(e^{i\pi\varepsilon}+e^{-i\pi\varepsilon})(e^{-\pi\xi}-e^{\pi\xi})}{e^{2i\pi\varepsilon}+e^{-2i\pi\varepsilon}-(e^{2\pi\xi}+e^{-2\pi\xi})}\right|=\left|\frac{2\cos{(\pi\varepsilon)}(e^{-\pi\xi}-e^{\pi\xi})}{2\cos{(2\pi\varepsilon)}-(e^{2\pi\xi}+e^{-2\pi\xi})}\right|$
$\leq\frac{2|e^{-\pi\xi}-e^{\pi\xi}|}{e^{2\pi\xi}+e^{-2\pi\xi}-2}=\frac{2|e^{-\pi\xi}-e^{\pi\xi}|}{|e^{\pi\xi}-e^{-\pi\xi}|^{2}}=\frac{2}{|e^{\pi\xi}-e^{-\pi\xi}|}.$
By (A.6), we see that $F_{\varepsilon}$ is uniformly bounded with respect to $\varepsilon$ for $|\xi|>r$; for instance, we have
$\left|F_{\varepsilon}(\xi)\right|\leq\frac{2}{e^{\pi r}-e^{-\pi r}}.$
Thus, Lebesgue’s dominated convergence theorem shows that
(A.7) $\lim_{\varepsilon\downarrow 0}\int_{|\xi|>r}F_{\varepsilon}(\xi)\phi(\xi)\,d\xi=\int_{|\xi|>r}\operatorname{cosech}{(\pi\xi)}\phi(\xi)\,d\xi.$
Furthermore, by (A.6), we see that
(A.8) $\lim_{\varepsilon\downarrow 0}\left|\int_{-r}^{r}F_{\varepsilon}(\xi)\left\{\phi(\xi)-\phi(0)\right\}d\xi\right|\leq\int_{-r}^{r}\frac{2e^{\pi\xi}|\xi|}{|e^{2\pi\xi}-1|}\left|\frac{\phi(\xi)-\phi(0)}{\xi}\right|d\xi\lesssim|r|\|\partial_{\xi}\phi\|_{L^{\infty}},$
where the implicit constant is independent of $\varepsilon$ and $r$. Putting (A.5), (A.7) and (A.8) together, we see that
(A.9) $\lim_{\varepsilon\downarrow 0}\int_{\mathbb{R}}F_{\varepsilon}(\xi)\phi(\xi)\,d\xi$ $=\lim_{r\downarrow 0}\lim_{\varepsilon\downarrow 0}\bigg\{\int_{|\xi|>r}F_{\varepsilon}(\xi)\phi(\xi)\,d\xi+\int_{-r}^{r}F_{\varepsilon}(\xi)\left\{\phi(\xi)-\phi(0)\right\}d\xi\bigg\}$
$=\lim_{r\downarrow 0}\int_{|\xi|>r}\operatorname{cosech}{(\pi\xi)}\phi(\xi)\,d\xi=\langle\textrm{p.v.}\operatorname{cosech}{(\pi\xi)},\phi\rangle.$
Thus, we have proven (A.4) and the lemma. ∎
Now, we are in a position to prove (1.1).
Proof of (1.1).
It is sufficient to prove that
(A.10) $\mathcal{F}[\textrm{p.v.}\coth{(\cdot)}](\xi)=-i\pi\,\textrm{p.v.}\coth{\Big(\frac{\pi\xi}{2}\Big)}.$
Indeed, (A.10) together with basic properties of Fourier transformations show that for any Schwartz function $f$ on $\mathbb{R}$,
$\mathcal{F}[T_{\delta}f](\xi)$ $=\frac{1}{2\delta}\mathcal{F}\left[\textrm{p.v.}\coth{\left(\frac{\pi\cdot}{2\delta}\right)}*f\right](\xi)=\frac{1}{2\delta}\mathcal{F}\left[\textrm{p.v.}\coth{\left(\frac{\pi\cdot}{2\delta}\right)}\right](\xi)\mathcal{F}[f](\xi)$
$=\frac{1}{\pi}\mathcal{F}\left[\textrm{p.v.}\coth{(\cdot)}\right]\Big(\frac{2\delta\xi}{\pi}\Big)\mathcal{F}[f](\xi)=-i\,\textrm{p.v.}\coth{\left(\delta\xi\right)}\mathcal{F}[f](\xi).$
Note that
$\frac{1}{e^{x}+1}-\frac{1}{e^{-x}+1}=\frac{e^{-\frac{x}{2}}-e^{\frac{x}{2}}}{e^{\frac{x}{2}}+e^{-\frac{x}{2}}}=-\tanh{\left(\frac{x}{2}\right)}.$
Hence, by Lemma A.1, we see that
(A.11) $\mathcal{F}\left[\tanh{\left(\frac{x}{2}\right)}\right]\left(\xi\right)=-\mathcal{F}\left[\frac{1}{e^{x}+1}-\frac{1}{e^{-x}+1}\right]\left(\xi\right)=-2i\pi\,\textrm{p.v.}\operatorname{cosech}{\left(\pi\xi\right)}.$
Furthermore, by (A.11), and $\mathcal{F}^{2}[f](x)=2\pi f(-x)$, we see that
(A.12) $\mathcal{F}\left[\textrm{p.v.}\operatorname{cosech}{\left(x\right)}\right](\xi)$ $=\frac{i}{2}\mathcal{F}\left[-2i\pi\,\textrm{p.v.}\operatorname{cosech}{\left(\pi x\right)}\right]\left(\pi\xi\right)$
$=\frac{i}{2}\mathcal{F}^{2}\left[\tanh{\left(\frac{x}{2}\right)}\right]\left(\pi\xi\right)=i\pi\tanh{\Big(\frac{-\pi\xi}{2}\Big)}=-i\pi\tanh{\Big(\frac{\pi\xi}{2}\Big)}.$
Note that
(A.13) $\coth{(x)}$ $=\frac{e^{x}+e^{-x}}{e^{x}-e^{-x}}=\frac{(e^{\frac{x}{2}}-e^{-\frac{x}{2}})^{2}+2}{e^{x}-e^{-x}}=\frac{e^{\frac{x}{2}}-e^{-\frac{x}{2}}}{e^{\frac{x}{2}}+e^{-\frac{x}{2}}}+\frac{2}{e^{x}-e^{-x}}$
$=\tanh{\left(\frac{x}{2}\right)}+\operatorname{cosech}{(x)}.$
Moreover, we see that
(A.14) $\operatorname{cosech}{\left(x\right)}+\coth{(x)}=\frac{2+e^{x}+e^{-x}}{e^{x}-e^{-x}}=\frac{(e^{\frac{x}{2}}+e^{-\frac{x}{2}})^{2}}{e^{x}-e^{-x}}=\coth{\left(\frac{x}{2}\right)}.$
Using (A.13) twice, (A.11), (A.12), and (A.14), we see that
(A.15) $\mathcal{F}\left[\textrm{p.v.}\coth{(x)}\right](\xi)$ $=\mathcal{F}\left[\tanh{\left(\frac{x}{2}\right)}\right](\xi)+\mathcal{F}\left[\textrm{p.v.}\operatorname{cosech}{(x)}\right](\xi)$
$=-2i\pi\,\textrm{p.v.}\operatorname{cosech}{\left(\pi\xi\right)}-i\pi\tanh{\Big(\frac{\pi\xi}{2}\Big)}$
$=-i\pi\,\textrm{p.v.}\operatorname{cosech}{\left(\pi\xi\right)}-i\pi\,\textrm{p.v.}\coth{(\pi\xi)}$
$=-i\pi\,\textrm{p.v.}\coth{\Big(\frac{\pi\xi}{2}\Big)}.$
Thus, we have proven (A.10); thus (1.1) is true. ∎
Appendix B Proof of Lemma 3.4 and (3.28)
Proof of Lemma 3.4.
(i) The first inequality follows immediately from the definition of $c_{j}[f]$. We also see that $c_{j\pm 1}[f]\leq 2^{\varepsilon}c_{j}[f]$, which implies the second inequality. For the last one, we easily see $\mathcal{E}^{2}(f)\lesssim\|\mathbf{c}[f]\|_{\ell^{2}}^{2}$, while the converse is shown by the (discrete) Young inequality.
(ii) Using the inequality $\big|\|\phi\|_{L^{2}}^{1/2}-\|\psi\|_{L^{2}}^{1/2}\big|\leq\|\phi-\psi\|_{L^{2}}^{1/2}$, we have
$|c_{j}[f]-c_{j}[g]|\leq\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\big\|P_{k}\partial_{x}(f-g)\big\|_{H^{1}}+2^{-\varepsilon j}\Big(\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}}^{\frac{1}{2}}\Big).$
The claim follows by Young’s inequality as before.
(iii) By Bernstein’s inequality, (i), and that $\varepsilon\in(0,1)$, we see that
$\|P_{\leq j}\partial_{x}f\|_{H^{2}}\lesssim\sum_{k=0}^{j}2^{k}\|P_{k}\partial_{x}f\|_{H^{1}}\leq\sum_{k=0}^{j}2^{k}2^{\varepsilon(j-k)}c_{j}[f]\lesssim 2^{j}c_{j}[f].$
Similarly,
$\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||P_{\leq j}f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}$ $\leq\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\big(\|P_{\leq j}f\|_{L^{\infty}}+\|f\|_{L^{\infty}}\big)\|P_{\geq j+1}f\|_{L^{2}}$
$\lesssim\|f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\sum_{k=j+1}^{\infty}\|P_{k}f\|_{L^{2}}^{2}$
$\lesssim\big(2^{\varepsilon j}c_{j}[f]\big)^{2}+\sum_{k=j+1}^{\infty}\big(2^{-2k}2^{\varepsilon(k-j)}c_{j}[f]\big)^{2}\lesssim 2^{2j}c_{j}[f]^{2}.$
The claim follows from these estimates. Thus, we have completed the proof of Lemma 3.4. ∎
Proof of (3.28).
By Bernstein’s inequality, the $\mathcal{Z}^{3}$ bound from (3.25), and Lemma 3.4 (iii), we see that
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\lesssim 2^{-k}\big(\|\partial_{x}u^{j}\|_{L^{\infty}_{T_{*}}H^{2}}+\|\partial_{x}u^{j-1}\|_{L^{\infty}_{T_{*}}H^{2}}\big)\lesssim 2^{-k}\big(\mathcal{E}^{3}(P_{\leq j}\phi)+\mathcal{E}^{3}(P_{\leq j-1}\phi)\big)^{\frac{1}{2}}$
$\lesssim 2^{-k+j}c_{j}[\phi].$
This gives (3.28) if $j\leq k$. When $k<j$, we use the difference bound (3.27) instead, to obtain
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}$ $\lesssim 2^{k}\|\partial_{x}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}\lesssim_{M_{0}}2^{k}d^{1}(P_{\leq j}\phi,P_{\leq j-1}\phi)$
$\lesssim 2^{k}\big(\|P_{j}\partial_{x}\phi\|_{L^{2}}+\|P_{j}\phi\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\|P_{j}\phi\|_{L^{2}}\big)$
$\lesssim 2^{k}\big(2^{-j}+2^{-\frac{3}{2}j}+M_{0}2^{-2j}\big)\|P_{j}\partial_{x}\phi\|_{H^{1}}\lesssim_{M_{0}}2^{k-j}c_{j}[\phi],$
as desired. ∎